### v1

# Heights of butterfly trees

Abstract
Binary search trees (BSTs) are fundamental data structures whose performance is largely governed by tree height. We introduce a block model for constructing BSTs by embedding internal BSTs into the nodes of an external BST—a structure motivated by parallel data architectures—corresponding to composite permutations formed via Kronecker or wreath products. Extending Devroye’s result that the height $h_{n}$ of a random BST satisfies $h_{n}/\log n\to c^{*}\approx 4.311$, we show that block BSTs with $nm$ nodes and fixed external size $m$ satisfy $h_{n,m}/\log n\to c^{*}+h_{m}$ in distribution. We then study butterfly trees: BSTs generated from permutations built using iterated such products. For simple butterfly trees (from iterated Kronecker products of $S_{2}$), we give a full distributional description showing polynomial height growth: $\mathbb{E}h_{n}^{\operatorname{B}}=\Theta(N^{\alpha})$ with $\alpha\approx 0.58496$. For nonsimple butterfly trees (from wreath products), we prove power-law bounds: $cN^{\alpha}\cdot(1+o(1))\leq\mathbb{E}h_{n}^{\operatorname{B}}\leq dN^{\beta}\cdot(1+o(1))$, with $\beta\approx 0.913189$.
1 Introduction
A binary tree is a simple hierarchical data structure, where every node can only have two immediate children (i.e., a left or right child), starting with a single root node. A binary search tree (BST) is a binary tree where the labeling of the nodes reflects the inherent ordering structure, where each left child and right child have, respectively, smaller and larger label (i.e., key) than the parent node. Nodes with no children are called leaves, and the depth of a node in the BST is the length of the path connecting the root to the node. The height $h$ of the BST is then the maximal depth among all nodes within the tree. The height is a fundamental property of BSTs as it directly determines the time complexity of standard operations such as search, insertion, and deletion, which are performed in $O(h)$ time. Binary search trees (BSTs) are widely used for managing ordered data in applications such as priority queues, file systems, and database indexing [12]. Recent work has also explored BST variants optimized for parallelism and cache efficiency, particularly in the context of concurrent data structures and large-scale data processing [7, 8, 3]. These developments motivate the study of alternative BST constructions that align more naturally with recursive data layouts and hierarchical memory architectures—directions that we pursue in this work.
A BST with $n$ nodes can be generated from an array ${\mathbf{x}}\in\mathbb{R}^{n}$ of distinct real numbers by iteratively applying the insertion operation using ${\mathbf{x}}$ as the vector of keys, starting with the empty tree. We will denote this tree as $\mathcal{T}({\mathbf{x}})$. The resulting underlying structure depends solely on the ordering of the keys, which can be derived from the corresponding permutation generated by the ordering. We define the depth of node $x_{k}$ as $\operatorname{depth}_{{\mathbf{x}}}(x_{k})$, which returns the number of edges traversed in $\mathcal{T}({\mathbf{x}})$ connecting the root $x_{1}$ to node $x_{k}$, with then the tree height $h(\mathcal{T}({\mathbf{x}}))=\max_{k}\operatorname{depth}_{{\mathbf{x}}}(x_{k})$. For example, the array ${\mathbf{x}}=[0.65,0.70,0.27,0.67,0.16,0.75]^{T}$ yields the permutation $\pi=352416\in S_{6}$, with the construction of $\mathcal{T}(\pi)$ outlined in Figure 1, that has height 2.
The behavior of BSTs generated from random permutations $\mathcal{T}(\pi)$ when $\pi\sim\operatorname{Unif}(S_{n})$ has been extensively studied (see [20, 15, 16, 39, 19, 17, 41]). A landmark result is Devroye’s theorem on the limiting height:
Theorem 1 ([15]).
If $\pi\sim\operatorname{Unif}(S_{n})$ and $h_{n}=h(\mathcal{T}(\pi))$, then $h_{n}/\log n$ converges in probability and in $L^{p}$ for $p\geq 1$ to $c^{*}\approx 4.311$ as $n\to\infty$, where $c^{*}$ denotes the unique solution to $x\log(2e/x)=1$ for $x\geq 2$.
More recent work has extended this result to broader classes of random permutations, including Mallows trees [2], record-biased trees [13] and permuton-sampled trees [14].
3
(a) Root 3: $\mathcal{T}(3)$
35
(b) Add 5: $\mathcal{T}(35)$
325
(c) Add 2: $\mathcal{T}(352)$
325146
(d) Adding 416, in sequence: $\mathcal{T}(352416)$
0.650.270.700.160.670.75
(e) Compare to $\mathcal{T}({\mathbf{x}})$
Figure 1: Step-by-step construction of $\mathcal{T}(352416)$ next to $\mathcal{T}({\mathbf{x}})$
We focus on block BSTs, a hierarchical structure in which internal BSTs are placed at each node of an external BST. This structure naturally arises in parallel implementations of BSTs (e.g., [11, 26]). We will consider the cases when the internal BSTs are identical as well as distinct (viz., they are independent and identically distribution (iid) in the random BST cases); these then correspond respectively to uniform permutations formed using Kronecker and wreath products (see Section 1.1 for definitions of each product), which we will denote $S_{n}\wr S_{m}$ and $S_{m}\otimes S_{n}$ (note the ordering of the $n,m$ terms differ, as on the wreath product the external structure is determined by the right action while in the Kronecker product the external structure is determined by the left action), that each comprise permutation subgroups $S_{m}\otimes S_{n}\subset S_{n}\wr S_{m}\subset S_{nm}$. The recursive construction differs from other studied recursive BST structures, such as seen in [38].
We are interested in how this construction compares to the uniform case on the BST with $nm$ nodes. A particular distinction when then studying the heights of block BSTs is how the internal BSTs are joined using the external BST. The heights are not additive across the internal BSTs, as the external BST edges join internal BSTs by traversing either the top-left or top-right edges (denoted $\ell_{n}$ and $r_{n}$ for a general BST with $n$ nodes) of the preceding internal BST before connecting to the root of the following internal BST. The top edges of the BST correspond explicitly to the left-to-right maxima paths (for the top-right edge) and the left-to-right minima paths (for the top-left edge), which are attained by taking a step up or down only when moving left to right from the first entry of the permutation $\pi$. For example, in Figure 1, 356 and 321 are the respective left-to-right maxima and minima paths seen within $\pi=352416$ which then manifest as the top-right and top-left edges of the corresponding BST $\mathcal{T}(\pi)$. (Note, for instance, that 4 does not appear within the left-to-right maxima path since 5 is encountered before 4 in $\pi$.)
For comparison, Kronecker and wreath products of permutations have recently been studied under the guise of the impact on the length of the longest increasing subsequence (LIS). The LIS problem has a very rich history of its own (see [42] for a thorough historical overview of the LIS problem). Among this history, we highlight the Vershik-Kerov theorem that $\operatorname{LIS}(\pi_{n})/\sqrt{n}$ converges in probability to 2 when $\pi\sim\operatorname{Unif}(S_{n})$. Very recently, Chatterjee and Diaconis [10] extended the Vershik-Kerov theorem to consider the wreath product of permutations; they established if both $m,n$ are growing sufficiently fast (viz., $n/\log^{4}m\to\infty$), then $\operatorname{LIS}(\pi)/\sqrt{nm}$ converges in probability to 4 when $\pi\sim\operatorname{Unif}(S_{n}\wr S_{m})$. A simple observation is the fact one wreath product is enough to increase the LIS from the uniform permutation case by a factor of 2 while maintaining the square root scaling. So one could then ask how do repeated wreath products impact the LIS, and can these change the scaling for the LIS? In [36], we resolve this question by showing that $n$-fold Kronecker or wreath products increase the power-law scaling of the LIS from $\sqrt{N}$ to at least $N^{\alpha}$, where $\alpha=\log_{2}(3/2)\approx 0.58496$ and $N=2^{n}$ is the length of the permutation. Here, we initiate a parallel investigation for the height of block BSTs.
Our first (motivational) contribution mimics the Chatterjee-Diaconis result when applied to the heights:
Theorem 2.
Let $\pi\sim\operatorname{Unif}(S_{n}\wr S_{m})$ or $\pi\sim\operatorname{Unif}(S_{m}\otimes S_{n})$, and $h_{n,m}=h(\mathcal{T}(\pi))$. If $m$ is fixed, then $h_{n,m}/\log(nm)$ converges in distribution to $c^{*}+h_{m}$ as $n\to\infty$.
Here, $c^{*}\approx 4.311$ is the exact same constant from Theorem 1, and $h_{m}$ denotes the (random) height of a BST with $m$ nodes. When $m=1$, this recovers Devroye’s original result.
This shows that even a single layer of block structure (via one wreath or Kronecker product) increases the limiting height, while preserving logarithmic scaling. We conjecture that when $n$ is fixed and $m\to\infty$, the properly rescaled height converges to $c^{*}H_{n}^{(1)}$, where $H_{n}^{(k)}=\sum_{j=1}^{n}1/{j^{k}}$ denotes the $n^{th}$ generalized harmonic number. For the case when both $n$ and $m$ grow, we conjecture the height scales as $\log n\log m$, exceeding the $\log(nm)$ scaling of a uniform $S_{nm}$ permutation.
The remainder of the paper explores how repeated Kronecker or wreath products affect the height of random BSTs, as well as the interplay between BST height and other permutation statistics such as the LIS, number of cycles, and left-to-right extrema. As an example, Figure 2 shows the BST for a permutation in $S_{30}$, highlighting LIS and LDS (longest decreasing sequence) paths. Unlike height, which follows downward paths constrained by the BST structure, the LIS can “jump” across branches—leading to polynomial, rather than logarithmic, scaling behavior.
3465178210911121314162017151821222523242627282930
Figure 2: $\mathcal{T}(\pi)$ for $\pi\in S_{30}$, with $h(\mathcal{T}(\pi))=6$ and particular LIS (from nodes 12 to 30) and LDS (from nodes 25 to 3) dotted paths are shown over the BST in red and green, respectively.
For this setting, we study the case of butterfly trees. Butterfly trees are explicitly the BSTs generated using butterfly permutations, which were introduced in [34, 36]. We let $\mathcal{T}_{n,s}^{\operatorname{B}}$ and $\mathcal{T}_{n}^{\operatorname{B}}$ denote, respectively, the simple and nonsimple butterfly trees with $N=2^{n}$ nodes. Butterfly permutations are permutations of length $N=2^{n}$, defined recursively via two constructions: simple butterfly permutations $\mathcal{B}_{n,s}$, built from $n$-fold Kronecker products (yielding the trees $\mathcal{T}_{n,s}^{\mathcal{B}}$), and nonsimple butterfly permutations $\mathcal{B}_{n}$, formed from $n$-fold wreath products (producing $\mathcal{T}_{n}^{\mathcal{B}}$). Both constructions begin from $S_{2}$. These permutations arise naturally as the pivoting patterns generated by Gaussian elimination with partial pivoting (GEPP) applied to random butterfly matrices—hence the name (see Section 1.1 for precise definitions). Structurally, the nonsimple butterfly permutations form a Sylow 2-subgroup of $S_{2^{n}}$, while the simple butterfly permutations form an abelian normal subgroup within them.
In addition to their link to numerical linear algebra, butterfly permutations and matrices have deeper connections to the study of automorphism groups of rooted trees, with implications across multiple areas including quantum information theory, random graph theory, and error-correcting codes (see Section 1.3 for more details).
Our main contributions are in an average-case analysis of the heights of butterfly trees with $N=2^{n}$ nodes, where we establish polynomial average height growth rather than the logarithmic growth for uniform permutations. For the case of simple butterfly trees, we can provide full nonasymptotic and asymptotic distributional descriptions of the heights:
Theorem 3.
Let $\pi\sim\operatorname{Unif}(\operatorname{B}_{n,s})$ and $h_{n}^{\operatorname{B}}=h(\mathcal{T}(\pi))$. Then $h_{n}^{\operatorname{B}}\sim 2^{X_{n}}+2^{n-X_{n}}-2$, where $X_{n}\sim\operatorname{Binom}(n,1/2)$, while $(\log_{2}h_{n}^{\operatorname{B}}-n/2)/(\sqrt{n}/2)$ converges in distribution to $|Z|$ as $n\to\infty$, where $Z\sim N(0,1)$. In particular, $\mathbb{E}h_{n}^{\operatorname{B}}=2(3/2)^{n}-2=cN^{\alpha}\cdot(1+o(1))$ with $c=2$ and $\alpha=\log_{2}(3/2)\approx 0.58496$.
The main ideas needed follow from establishing $h_{n}^{\operatorname{B}}=\ell_{n}^{\operatorname{B}}+r_{n}^{\operatorname{B}}$, along with then showing $\ell_{n}^{\operatorname{B}}=\operatorname{LIS}(\pi)-1$ and $r_{n}^{\operatorname{B}}=\operatorname{LDS}(\pi)-1$ for random simple butterfly permutations $\pi\sim\operatorname{Unif}(\operatorname{B}_{n,s})$, which is combined with the results $\log_{2}\operatorname{LIS}(\pi)\sim\operatorname{Binom}(n,1/2)$ and $\operatorname{LIS}(\pi)\cdot\operatorname{LDS}(\pi)=N$ that were each established in [36]. This marks a clear distinction from the uniform permutation of length $n$ case: $h_{n}$, $\ell_{n}$, and $r_{n}$ converge in probability, respectively, to $c^{*}\approx 4.311,1,1$ when scaled by $\log n$, so typically $h_{n}>2(\ell_{n}+r_{n})$, while also $\operatorname{LIS}(\pi)$ is of a polynomial order $N^{1/2}$, much larger than the logarithmic scalings observed for the other statistics. For butterfly trees, then the heights and LIS on average are now on the exact same order. We also highlight additional connections of the heights to degree distributions of vertices on the comparability graph for the Boolean lattice with $n$-element ground set.
Our next contribution is in regard to $h_{n}^{\operatorname{B}}$ for the case of nonsimple butterfly trees. In this setting, we no longer can provide full distributional descriptions. Instead, we focus on establishing lower and upper power-law bounds on the first moment $\mathbb{E}h_{n}^{\operatorname{B}}$.
Theorem 4.
Let $\pi\sim\operatorname{Unif}(\operatorname{B}_{n})$ and $h_{n}^{\operatorname{B}}=h(\mathcal{T}(\pi))$. Then $cN^{\alpha}\cdot(1+o(1))\leq\mathbb{E}h_{n}^{\operatorname{B}}\leq dN^{\beta}\cdot(1+o(1))$ for $\alpha\approx 0.58496$, $\beta\approx 0.913189$, $c=2$, and $d\approx 2.60958$.
The lower bound matches exactly the first moment for the simple case in Theorem 3. A similar behavior was observed for the $\operatorname{LIS}$ of nonsimple butterfly permutations in [36], where again only power-law bounds were attained for the first moment $\mathbb{E}\operatorname{LIS}(\pi)$. The main tools for Theorem 4 involve leveraging and iterating recursive relationships between the first two moments of $h_{n}^{\operatorname{B}}$, along with establishing additional properties for the top-left and top-right edge lengths—$\ell_{n}^{\operatorname{B}}$ and $r_{n}^{\operatorname{B}}$—for butterfly trees. Unlike in the simple case, where $r_{n}^{\operatorname{B}}$ aligned precisely with the $\operatorname{LIS}$ of the corresponding butterfly permutations, we instead show that $r_{n}^{\operatorname{B}}$ now aligns with the number of cycles for this permutation. The number of cycles of butterfly permutations was more thoroughly studied in [36] than the LIS question in the nonsimple setting, as full nonasymptotic and asymptotic results were attainable for this statistic (that have behavior determined by moments that satisfy further recursive relations). It is also less clear how $h_{n}^{\operatorname{B}}$ compares to $\ell_{n}^{\operatorname{B}}+r_{n}^{\operatorname{B}}$, other than $h_{n}^{\operatorname{B}}$ should (at least asymptotically) be larger than $\ell_{n}^{\operatorname{B}}+r_{n}^{\operatorname{B}}$ that have mean still proportional to the above lower bound $N^{\alpha}$.
Future directions and extensions.
A natural next step is to refine our understanding of the nonsimple butterfly case by seeking an exact expression—or at least tighter asymptotic characterization—for the average height of random butterfly trees, if not a full distributional description. In contrast to the simple case, where a precise distributional description is possible, closing this gap for the average heights in the nonsimple setting remains an open challenge. Additionally, establishing sharper asymptotic results for general block BSTs with non-fixed external block size presents another important direction for future work, as well as exploring similar directions for $p$-nary butterfly trees.
Beyond uniform, it would be valuable to explore butterfly and block permutation models with additional structural biases. One promising direction is to introduce Mallows-type distributions over butterfly or Kronecker-block permutations, which incorporate a bias parameter $q>0$ to favor or penalize certain block structures. This would generalize our current results in the same way that the Mallows model generalizes the uniform case for permutations. It would also further extend classical results such as those of Devroye (cf. Theorem 1), whose analysis of the expected height of uniform BSTs is recovered in the 1-block case of our framework or the $q=1$ balanced Mallows model. These extensions could lead to a broader theory of BST height behavior under structured and biased permutation models.
Outline.
The remaining paper is organized as follows. The remainder of Section 1 reviews relevant background on uniform permutations and introduces the formal definitions of butterfly permutations. The subsequent sections focus on the proofs of our main results and related discussion. Section 2 addresses Theorem 2 regarding block BSTs, while Section 3 covers both Theorems 3 and 4 for butterfly trees.
1.1 Preliminaries and notation
Let ${\mathbf{e}}_{i}\in\mathbb{R}^{n}$ denote the standard basis vectors, with 1 in the $i^{th}$ component and 0 elsewhere. For matrices $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{m\times m}$, let $A\otimes B\in\mathbb{R}^{nm\times nm}$ denote the Kronecker product formed by inserting $B$ into each element of $A$, as
$A\otimes B=\begin{bmatrix}A_{11}B&\cdots&A_{1n}B\\ \vdots&\ddots&\vdots\\ A_{n1}B&\cdots&A_{nn}B\end{bmatrix}.$
The Kronecker product satisfies the mixed-product property, $(A\otimes B)(C\otimes D)=(AC)\otimes(BD)$, that can be read as the product of Kronecker products is the Kronecker product of products (assuming compatible matrix dimensions). We will also write $A\oplus B$ for the block diagonal matrix with corresponding diagonal blocks $A_{1},A_{2}$. Let $S_{n}$ denote the symmetric group that consists of the permutations $\pi$ of a set of $n$ objects, and let $\mathcal{P}_{n}=\{P_{\pi}:\pi\in S_{n}\}$ denote the permutation matrices defined by $P_{\pi}{\mathbf{e}}_{i}={\mathbf{e}}_{\pi(i)}$ (i.e., the left-regular representation of the action of $S_{n}$ on the standard basis vectors). For permutations $\pi\in S_{n},\rho\in S_{m}$, let $\pi\otimes\rho\in S_{n}\otimes S_{m}\subset S_{nm}$ denote the corresponding permutation attained by $P_{\pi\otimes\rho}=P_{\pi}\otimes P_{\rho}$. Similarly, let $\pi=\pi_{1}\oplus\pi_{2}$ denote the corresponding permutation with permutation matrices $P_{\pi}=P_{\pi_{1}}\oplus P_{\pi_{2}}$, while $\pi=\pi_{1}\ominus\pi_{2}$ similarly denotes permutations where $P_{\pi}=P_{\pi_{1}}\ominus P_{\pi_{2}}$, where
$A\ominus B=\begin{bmatrix}{\mathbf{0}}&B\\ A&{\mathbf{0}}\end{bmatrix}.$
For subgroups $G\subset S_{n}$, $H\subset S_{m}$, let $G\wr H=G^{m}\rtimes H$ denote the wreath product formed by the action of $\rho\in H$ on the indices of an $m$-vector with components taken from copies of $G$ via
$\rho\cdot(x_{1},x_{2},\ldots,x_{m})=(x_{\rho(1)},x_{\rho(2)},\ldots,x_{\rho(m)}).$
Note further then $S_{m}\otimes S_{n}\subset S_{n}\wr S_{m}$ as the Kronecker products comprise the subcases of the wreath products when the internal blocks are identical.
For a sequence of random variables $X_{1},X_{2},\ldots$ and a random variable $Y$ (defined on the same sample space), we write $X_{n}\to_{p}Y$ to denote the sequence $X_{n}$ converges in probability to $Y$, $X_{n}\to_{d}Y$ to denote the sequence $X_{n}$ converges in distribution to $Y$, and $X_{n}\to_{L^{p}}Y$ to denote the sequence $X_{n}$ converges in $L^{p}$ to $Y$. We also recall standard results that will be utilized several times throughout, such as the equivalence of convergence in probability and distribution when the limit is constant, as well as Slutsky’s theorem:
Theorem 5 (Slutsky’s theorem, [43]).
If $X_{n},Y_{n}$ are sequences of random variables such that $X_{n}\to_{d}X$ and $Y_{n}\to_{p}\nu$ for a random variable $X$ and a constant $\nu$, then $X_{n}+Y_{n}\to_{d}X+\nu$ and $X_{n}Y_{n}\to_{d}\nu X$.
Additionally, we will make use of the Subgroup Algorithm due to Diaconis and Shashahani [18]:
Theorem 6 (Subgroup algorithm, [18]).
If $G$ is a compact group, $H$ is a closed subgroup of $G$, and $G/H$ is the set of left-cosets of $H$ in $G$, then $x\sim\operatorname{Unif}(G/H)$ and $y\sim\operatorname{Unif}(H)$ are independent if and only if $xy\sim\operatorname{Unif}(G)$.
This result allows for uniform sampling from $G$ by independently sampling a uniform coset representative and a uniform element from the subgroup $H$.
1.2 Random BSTs
Here we review briefly the background for the heights of random BSTs along with the lengths of the top-left and -right edges of the BST. Random BSTs are generated using uniform permutations, and have been the primary setting for the initial analysis of BST data structures. The height, $h_{n}$, for a BST with $n$ nodes is a fundamental statistic that governs the performance of common operations on the BST, such as search, insertion, and deletion, that each run in $O(h_{n})$ time. A trivial lower bound of $h_{n}\geq\lfloor\log_{2}n\rfloor$ is attained by the completely balanced BST (where each parent node has exactly two children when $n=2^{m+1}$), while a worst-case of $h_{n}\leq n-1$ is attained by the identity permutation (that consists only of a single right-edge path connecting the root 1 to the leaf node $n$). Further balancing mechanisms (e.g., AVL or Red-Black trees) can enforce $O(\log n)$ height but adds further complexity to the implementation that proves to be unnecessary on average (see below).
The average-case behavior of BSTs constructed under random insertion of $n$ distinct elements reveals insights into the practical performance of BSTs, as randomness tends to produce efficient tree structures. Devroye’s seminal result (see Theorem 1) establishes that for a BST built from $n$ random insertions, the height is $O(\log n)$. (Later work further establishes the variance is bounded [17, 19].) Hence, the average-case efficiency of BSTs aligns closely with the best-case $O(\log n)$ performance, differing only by a constant factor. Specifically, the height constant from the best-case balanced model is $1/\log 2\approx 1.4427$, compared to the exact constant $c^{*}\approx 4.31107$ for the average-case model confirmed by Devroye. Devroye’s result built on previous work of Pittel [37], who established $h_{n}/\log n$ converged almost surely to a positive constant, establishing bounds on this constant between 3.58 and 4.32. To recover the explicit constant $c^{*}$ required heavy machinery under the hood, using a correspondence between random BSTs and branching processes along with then the Hammersley-Kingman-Biggins theorem [6, 23, 27] that provides a law of large numbers result for the minimum on a branching random walk (necessary to match the lower bound to $c^{*}$).
Heights of more general random BST models have also been explored more recently, including the Mallows trees [2], record-biased trees [13], and permuton sampled trees [14]. Of these, we further highlight Mallows trees, which are generated using Mallows permutations [29], that bias permutations closer or farther from the identity permutation through a scaling $q^{\operatorname{inv}(\pi)}$ for $\operatorname{inv}$ returning the number of inversions of a given permutation with a weight $q\in(0,\infty)$; where $q=1$ aligns with the uniform permutation setting (see also [4, 30, 5, 22, 24, 25, 45, 46]). When $q\in(0,1)$, then Mallows permutations are biased more toward the identity permutation, and so the limiting behavior relies more heavily on the height of the right subtree in the permutation. Hence, in particular, these rely also on the length of the top-right edge of the BST, connecting the root to the node $n$. This corresponds to the left-to-right maxima path within the permutation, as we will review now. (Note this applies exactly analogously to the left-to-right minima path, that corresponds to the top-left edge $\ell_{n}$ of the corresponding BST.)
The left-to-right maxima path in a permutation $\pi\in S_{n}$ is attained by recording the maximum in the left subsequences of $\pi$ of length $k$, with $k=1,2,\ldots,n$. This path stops as soon as $n$ is encountered, since $n$ is necessarily the overall maximum. This is seen in Figure 3 as the permutation 216534 has left-subsequences 2, 21, 216 until the maximal 6 entry is encountered, so the left-to-right maxima subsequence consists of just 26. This subsequence explicitly materializes in the corresponding BST for $\pi\in S_{n}$ as this is determined by the nodes visited in the path connecting the root $\pi(1)$ to $n$ (e.g., in Figure 3, the top-right edge consists of the root 2 and 6). Hence, if $r_{n}$ is the length of the top-right edge of the corresponding BST $\mathcal{T}(\pi)$, then the length of the left-to-right maxima path inside $\pi$ is $r_{n}+1$.
For a uniform permutation, the distribution of the length of the left-to-right maxima has been understood since the 1960s, due to early work by Rényi [40]. In particular, one can show then $r_{n}+1\sim\Upsilon_{n}$, where $\Upsilon_{n}$ denotes the Stirling-1 distribution with probability mass function (pmf) given by
$\mathbb{P}(\Upsilon_{n}=k)=\frac{|s(n,k)|}{n!}$ (1)
for $k=1,2,\ldots,n$, where $s(n,k)$ denotes the Stirling numbers of the first kind. More traditionally, $\Upsilon_{n}$ models the number of cycles $C(\pi)$ in the disjoint cycle decomposition of a uniform permutation $\pi$; this follows as $|s(n,k)|$ can be defined to provide the number of permutations in $S_{n}$ with $k$ cycles in their disjoint cycle decomposition. (It follows then directly that (1) defines a bona fide pmf since then $\sum_{k=1}^{n}|s(n,k)|=|S_{n}|=n!$.) The Stirling numbers of the first kind can be computed explicitly using the recursion
$|s(n+1,k)|=n|s(n,k)|+|s(n,k-1)|.$ (2)
A standard interpretation for this recursion is building up a permutation from $\pi\in S_{n}$ to a permutation of $S_{n+1}$ by considering where $n+1$ is placed between the elements from $\pi$, with the cases being split between $n+1$ being appended to the end of $\pi$ versus being placed at any earlier space. To verify then $r_{n}+1\sim\Upsilon_{n}$, it suffices to show the left-to-right maxima can similarly decompose the permutations of length $n$ into sets of size $|s(n,k)|$ where the permutation has left-to-right maxima path of length $k$, and so also satisfy (2). This follows exactly as in the cycle case, but now following how 1 can be distributed within the permutation of length $n+1$: the placement of 1 only impacts the length of the left-to-right maxima path if $\pi(1)=1$, as any other placement means 1 will never be visited on the maxima path. Standard results establish $\mathbb{E}\Upsilon_{n}=H_{n}^{(1)}$ and $\operatorname{Var}\Upsilon_{n}=H_{n}^{(1)}-H_{n}^{(2)}$. Since $H_{n}^{(1)}=\log n+\gamma+o(1)$ for $\gamma\approx 0.577$ the Euler-Mascheroni constant and $H_{n}^{(2)}=\frac{\pi^{2}}{6}+o(1)$, along with the fact $h_{n}$ dominates $r_{n}+1$, then standard probability tools (e.g., Chebyshev’s inequality, uniform integrability) yield:
Corollary 1.
$\Upsilon_{n}/\log n$ converges in probability and in $L^{p}$ to 1 for all $p\geq 1$.
The left-to-right maxima and minima paths play a more pronounced role in the study of the heights of block BSTs, as will be explored in Section 2.
1.3 Butterfly permutations
A large focus of our paper will be on butterfly trees, which are BSTs derived from two classes of permutations of length $N=2^{n}$ called butterfly permutations. These are:
  • the simple butterfly permutations, $\operatorname{B}_{n,s}=S_{2}\otimes\cdots\otimes S_{2}=S_{2}^{\otimes n}$, an $n$-fold iterated Kronecker product of $S_{2}$, and
  • the nonsimple butterfly permutations, $\operatorname{B}_{n}=S_{2}\wr\cdots\wr S_{2}=S_{2}^{\wr n}$, an $n$-fold iterated wreath product of $S_{2}$.
Both comprise subgroups of $S_{N}$, with $\operatorname{B}_{n}$ a 2-Sylow subgroup of $S_{N}$ while $\operatorname{B}_{n,s}$ an abelian normal subgroup of $\operatorname{B}_{n}$. The term butterfly permutations originates from [34, 36], where they arise as the permutation matrix factors resulting from applying Gaussian elimination with partial pivoting (GEPP) to scalar butterfly matrices—specifically, when GEPP is applied to uniformly sampled butterfly matrices, the resulting permutations are uniformly sampled from $\operatorname{B}_{n}$ and $\operatorname{B}_{n.s}$.
The associated butterfly matrices are recursively defined block-structured matrices generated using classical Lie group elements, especially $\operatorname{SO}(2)$ (or more generally $\operatorname{U}(2)$; see [32]). In particular, the simple butterfly matrices are of the form $\bigotimes_{j=1}^{n}\operatorname{SO}(2)$, which itself comprises an abelian subgroup of $\operatorname{SO}(N)$, and so has an induced Haar measure. Nonsimple butterfly matrices are formed recursively, starting with $\operatorname{SO}(2)$, and the order $N=2^{n}$ butterfly matrices are formed by taking the product
$\begin{bmatrix}C&S\\ -S&C\end{bmatrix}\begin{bmatrix}A_{1}&{\mathbf{0}}\\ {\mathbf{0}}&A_{2}\end{bmatrix}=\begin{bmatrix}CA_{1}&SA_{2}\\ -SA_{1}&CA_{2}\end{bmatrix}$
where $A_{1},A_{2}$ are order $N/2$ butterfly matrices, and $C,S$ are symmetric commuting matrices that satisfy the Pythagorean matrix identity $S^{2}+C^{2}={\mathbf{I}}$. The scalar butterfly matrices assume the $(C,S)$ pair used at each recursive step are the scalar matrices $(C,S)=(\cos\theta,\sin\theta){\mathbf{I}}$ for some angle $\theta$; in this case, the above form condenses to
$(R_{\theta}\otimes{\mathbf{I}})(A_{1}\oplus A_{2})$
for $R_{\theta}$ the (clockwise) rotation matrix of angle $\theta$. The simple butterfly matrices enforce the additional constraint that $A_{1}=A_{2}$ at each intermediate recursive step, that allows the further collapsing of the above form into an iterated Kronecker product. The mixed-product property then derives the necessary Kronecker product form of the permutation matrix factors of simple scalar butterfly matrices. The nonsimple case is less direct (see [36, Theorem 5]); for instance, the nonsimple butterfly matrices are not closed under multiplication but the associated GEPP permutation matrix factors still comprise a group.
The sizes of each butterfly permutation group are
$|B_{n,s}|=2^{n}=N,\qquad|B_{n}|=2^{2^{n}-1}=2^{N-1}.$
This confirms $\operatorname{B}_{n}$ as a 2-Sylow subgroup of $S_{2^{n}}$, as it matches the number of 2-factors in $|S_{2^{n}}|=(2^{n})!$. Similar constructions generalize to $p$-Sylow subgroups of $S_{p^{n}}$ for other primes $p$ (see [36]).
Butterfly matrices have been widely studied in numerical linear algebra, especially for their efficiency and structure in fast solvers [44, 35, 34, 33, 28]. Their recursive structure, when constructed from unitary groups, makes butterfly matrices natural models for quantum entanglement [31]. This perspective extends to the associated butterfly permutations, whose group structure, low circuit complexity, and recursive symmetry make them well-suited as efficient data structures for quantum information processing.
These butterfly permutation groups also have deep connections to the automorphism group of the infinite rooted $p$-nary tree. For instance, Abért and Virág [1] resolve a classical question of Turán regarding the typical order of an element from such a $p$-Sylow subgroup utilizing this correspondence. This perspective connects butterfly groups to key questions in random graph theory, such as the girth of random Cayley graphs [21] and optimal mixing in Markov chains [9]. In this paper, we focus on the binary case $p=2$.
2 Heights of block BSTs
We first study the height behavior of block BSTs—hierarchical structures formed by replacing each node of an external BST with an internal BST. These arise naturally in settings involving the parallelization of BST operations, particularly where permutations define structural evolution. We now focus on two combinatorially natural ways of generating these structures: via the wreath product and the Kronecker product of permutations.
21
(a) $\mathcal{T}(12)$
21
(b) $\mathcal{T}(21)$
213
(c) $\mathcal{T}(132)$
652143
(d) $\mathcal{T}(216534)$
Figure 3: $\mathcal{T}(216534)$ formed using $\mathcal{T}(132)$ with nodes replaced by copies of $\mathcal{T}(12)$ and $\mathcal{T}(21)$.
We first consider the question of the heights of corresponding BSTs generated using permutations sampled from $S_{n}\wr S_{m}$ and $S_{m}\otimes S_{n}$. We will refer to the $S_{m}$ as the external permutation model for this product, and the $S_{n}$ as corresponding to the internal permutation models. Note these labels align, in particular, with how the corresponding permutation matrix for such a product could be formed: the $S_{m}$ corresponding permutation matrix determines the locations of the $n\times n$ blocks within the $nm\times nm$ permutation matrix, and each $n\times n$ block then consists of a particular permutation matrix corresponding to the $S_{n}$ factor. For example, consider $132\in S_{3}$, along with the permutations $21,12,21\in S_{2}$, and then $(21\oplus 12\oplus 21)\rtimes 132=(21|34|65)\rtimes 132=216534\in S_{2}\wr S_{3}\subset S_{6}$ has corresponding permutation matrix $P_{216534}$ formed by placing copies of $P_{12}$ and $P_{21}$ within $P_{132}$, as seen here:
$P_{216534}=\begin{bmatrix}\mathbf{0}&\mathbf{1}&0&0&0&0\\ \mathbf{1}&\mathbf{0}&0&0&0&0\\ 0&0&0&0&\mathbf{1}&\mathbf{0}\\ 0&0&0&0&\mathbf{0}&\mathbf{1}\\ 0&0&\mathbf{0}&\mathbf{1}&0&0\\ 0&0&\mathbf{1}&\mathbf{0}&0&0\end{bmatrix}=\begin{bmatrix}P_{21}&{\mathbf{0}}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{0}}&P_{12}\\ {\mathbf{0}}&P_{21}&{\mathbf{0}}\end{bmatrix}=(P_{21}\oplus P_{12}\oplus P_{21})\left(P_{132}\otimes{\mathbf{I}}_{2}\right)$
This block structure materializes then directly within the BST for $\mathcal{T}(216534)$ as this can be realized as the external BST for $\mathcal{T}(132)$ being updated by replacing each node with copies of $\mathcal{T}(12)$ or $\mathcal{T}(21)$ (with keys shifted according to the corresponding labels from the external BST), as seen in Figure 3. In particular, we note that while the nodes of the external BST are then replaced by internal BST copies, the edges from the external BST that glue together each of the node BSTs connect along the top-left or top-right edges of the parent external node to the root of the child external node (e.g., in Figure 3, the external node 3 is the parent for external node 2, so the corresponding internal BST $\mathcal{T}(65)$ connects along the top-left edge (that ends at 5) to the root of $\mathcal{T}(34)$; so the external edge connects 5 directly to 3 in the overall block BST). Figure 4 shows a particular instance of a block BST formed using Kronecker products instead, where repeated copies of $\mathcal{T}(2143)$ are then inserted into the node locations of $\mathcal{T}(\rho)$ with $\rho\in S_{2}$ to then form $\mathcal{T}(\rho\otimes 2143)$.
2143
(a) $\mathcal{T}(2143)$
65827143
(b) $\mathcal{T}(21436587)=T(12\otimes 2143)$
65827143
(c) $\mathcal{T}(65872143)=T(21\otimes 2143)$
Figure 4: $\mathcal{T}(12\otimes 2143)$ and $\mathcal{T}(21\otimes 2143)$ formed from gluing two copies of $\mathcal{T}(2143)$ together with an extra edge (in blue) connecting the top-left edges or top-right paths of each parent BST to the root of the child BST.
As we are considering the heights of these block BST models, we note these heights do not simply add, due to the dependence on the paths used to reach the leaves across internal and external blocks. So we need to have a clear understanding of the top-left and top-right traversals taken within each internal BST to get to the leaves of the overall block BST. For example, in Figure 3, we have a height of 4, that is attained by traversing the 2 external edges along with 2 internal edges. Let $\ell(i),r(i),h(i)$ denote the corresponding top-left edge, top-right edge, and height for the internal BST found in the external node $i$, then we realize $h:=h(\mathcal{T}(216534))=4$ as $h=r(1)+1+\ell(3)+1+h(2)$ (noting further $r(1)=0$ as the right sub-tree of node 2 is empty for $\mathcal{T}(21)$). We thus emphasize the more explicit reliance on understanding the behavior of the top-left and top-right edges for the height of the overall block BST, which correspond to the paths from the root to 1 and $n$, respectively.
We are interested in the asymptotic heights of random block BSTs generated using products of permutations of sizes $n,m$ where $nm$ grows. We first focus on Theorem 2, that studies the heights of $\mathcal{T}(\pi)$ where $\pi\sim\operatorname{Unif}(S_{n}\wr S_{m})$ or $\pi\sim\operatorname{Unif}(S_{m}\otimes S_{n})$ in the case when $m$ is fixed and $n$ grows unbounded. This is a generalization of Theorem 1, that is thus subsumed by the case $m=1$.
Note $S_{m}\otimes S_{n}\subset S_{n}\wr S_{m}$ naturally, as the left hand side comprises the subcases when the internal blocks are identical inside a wreath product structure (and so the semi-direct product action is trivial). Let $\pi\in S_{n}\wr S_{m}$. We can further identify the coordinates of $\pi$ by the $m$-node block locations of each set of coordinates, by writing $\pi=(\pi_{\rho(1)}|\pi_{\rho(2)}|\cdots|\pi_{\rho(m)})$, where $\rho\in S_{m}$ determines the external block structure of $P_{\pi}$, while $\pi_{j}$ are each shifted copies of permutations from $S_{n}$ (i.e., $\pi_{j}-\min(\pi_{j})+1\in S_{n}$) and thus each comprise the internal permutation matrices found within the blocks of $P_{\pi}$. Hence, $\mathcal{T}(\pi)$ can be equivalently generated by replacing the nodes of $\mathcal{T}(\rho)$ with copies of $\mathcal{T}(\pi_{j})$ that are connecting from parent to child blocks from $\mathcal{T}(\rho)$ along either the top-left or top-right edge of the parent block to the root of the child block (cf. Figures 3 and 4). In particular, note how the structure of $\mathcal{T}(\pi_{j})$ only cares about the order of the nodes within $\pi_{j}$ and so is agnostic of shifts of indices; hence $\mathcal{T}(\pi_{j})$ and $\mathcal{T}(\pi_{j}-\min(\pi_{j})+1)$ only differ by shifting each label by $-\min(\pi_{j})+1$ (cf. $\mathcal{T}(2143)$ and $\mathcal{T}(6587)$ within Figure 4). We are now prepared to prove Theorem 2.
Proof of Theorem 2.
We will only include the proof for the wreath product as the Kronecker product case would follow from the exact same argument adapted to consider identical internal blocks. By the subgroup algorithm (Theorem 6), independently sampling an element from the outer group $S_{n}$ and each component of the inner group $S_{m}$ yields a uniform sample from the full wreath product $S_{n}\wr S_{m}$. That is, if $\pi=(\pi_{\rho(1)}|\pi_{\rho(2)}|\cdots|\pi_{\rho(m)})\sim\operatorname{Unif}(S_{n}\wr S_{m})$ or $\pi\sim\operatorname{Unif}(S_{m}\otimes S_{n})$ with induced block permutations $\rho\in S_{m}$ and $\pi_{j}^{\prime}=\pi_{j}-\min(\pi_{j})+1\in S_{n}$ (i.e., a relabeling of the indices), then $\rho\sim\operatorname{Unif}(S_{m})$ is independent of each subblock induced permutation $\pi_{j}^{\prime}\sim\operatorname{Unif}(S_{n})$.
To compute the height of the block BST formed using $\pi\in S_{n}\wr S_{m}$, we need to find the maximal depth of any node within each block BST. Hence, we can first focus on finding the maximal depths within each $\pi_{j}$, and then finding the maximum among all of these computed maxima. In the case when $m$ is fixed and $n$ grows unbounded (as needed for Theorem 2), this two step process is completely tractable. The first step we will follow shows the depth of the node within a block in the overall tree is determined by the depth of the node inside of the internal permutation plus a shared path from the overall root to the root of the internal node block along the parental heritage paths consisting of top-left or top-right traversals from parent block to subsequent child block root. The next lemma formalizes how a node’s depth in the overall block BST decomposes into contributions from internal and external traversals.
Lemma 1.
Let $\pi=(\pi_{\rho(1)}|\cdots|\pi_{\rho(m)})\in S_{n}\wr S_{m}$, where $\rho\in S_{m}$ and $\pi^{\prime}_{j}=\pi_{\rho(j)}-\min(\pi_{\rho(j)})+1\in S_{n}$ for each $j$. For node $v$ in $\mathcal{T}(\pi_{\rho(j)})\subset\mathcal{T}(\pi)$, let $x=v-\min(\pi_{\rho(j)})+1$, and ${\mathbf{y}}=(y_{1},y_{2},\ldots,y_{k})$ the subsequence of $\rho$ that denotes the node sequence in the shortest path in $\mathcal{T}(\rho)$ connecting the root $\rho(1)=y_{1}$ to node $\rho(j)=y_{k}$. Let $\ell_{n}^{(j)}=\ell(\pi_{j})$ and $r_{n}^{(j)}=r(\pi_{j})$ denote the top-left and top-right edge lengths of $\mathcal{T}(\pi_{j})$, and define
$G(x,y,u,v)=\left\{\begin{array}[]{ll}u,&x>y,\\ v,&x<y.\end{array}\right.$
Then
$\operatorname{depth}_{\pi}(v)=\operatorname{depth}_{\pi^{\prime}_{j}}(x)+\sum_{k=1}^{\operatorname{depth}_{\rho}(\rho(j))}G(y_{k},y_{k+1},\ell_{n}^{({y_{k}})}+1,r_{n}^{({y_{k}})}+1).$
Lemma 1 captures the particular path within $\mathcal{T}(\pi)$ connecting the root $\pi(1)$ to the node $v$ is attained in an explicit form by the path within the corresponding $n$-node BST $\mathcal{T}(\pi_{\rho(j)})$ along with the path connecting the blocks that thus travels through either the top-left or top-right edges of the preceding parent $m$-node blocks, determined by the particular path in $T(\rho)$ connecting the root $\rho(1)$ to the corresponding $m$-node block location $\rho(j)$. Hence, the proof of Lemma 1 is immediate from construction. Note in particular the external $m$-node path is shared for all nodes within the same $m$-block. It follows then the maximal depth for any node within the same $m$-block $\rho(j)$ is determined by the height of $\pi_{j}^{\prime}$ and the depth of $\rho(j)$:
Corollary 2.
Let $\pi,\rho,\pi_{j}^{\prime}$, $x$, ${\mathbf{y}}$ be as in Lemma 1. Then
$\max_{v\in\mathcal{T}(\pi_{\rho(j)})}\operatorname{depth}_{\pi}(v)=h(\mathcal{T}(\pi_{j}^{\prime}))+\sum_{k=1}^{\operatorname{depth}_{\rho}(\rho(j))}G(y_{k},y_{k+1},\ell_{n}^{({y_{k}})}+1,r_{n}^{({y_{k}})}+1).$ (3)
Write $X_{n,j,k}:=G(y_{k},y_{k+1},\ell_{n}^{(y_{k})}+1,r_{n}^{(y_{k})}+1)$, where we note additionally $X_{n,j,k}\sim\Upsilon_{n}$ iid for each $n,j,k$. Conditioning on $\rho$ while then using Corollary 1 and $\Upsilon_{n}\sim\ell_{n}^{(y_{k})}+1\sim r_{n}^{(y_{k})}+1$, then we have $X_{n,j,k}/\log n$ converges to 1 in probability and $L^{p}$ for all $p\geq 1$ as $n\to\infty$. Similarly, $h(\mathcal{T}(\pi_{j}^{\prime}))/\log n$ converges to $c^{*}$ in probability and $L^{p}$ for all $p\geq 1$, so that the scaled right hand side of (3) converges to
$c^{*}+\operatorname{depth}_{\rho}(\rho(j))$
in probability (still conditioning on $\rho$) by Slutsky’s theorem (Theorem 5). Now let ${\mathbf{z}}\in\mathbb{R}^{m}$ be the vector with components $z_{j}=\max_{v\in\mathcal{T}(\pi_{\rho(j)})}\operatorname{depth}_{\pi}(v)$ from which it follows
$h(\mathcal{T}(\pi))=\max_{j}z_{j}.$
In particular, $h(\mathcal{T}(\pi))$ is a continuous map of ${\mathbf{z}}\in\mathbb{R}^{m}$.
If $m$ is fixed, then $h(\mathcal{T}(\pi_{j}^{\prime}))/\log n\to_{p}c^{*}$ while $X_{n,j,k}/\log n\to_{p}1$ (by Theorem 1 and Corollary 1). Hence, using the equivalence of limits in probability and in distribution when the limiting term is constant along with the independence of $\pi_{j}^{\prime}$ and $\rho$, we see for fixed $j\leq m$ then
$\lim_{n\to\infty}\mathbb{P}\left(\sum_{k=1}^{\operatorname{depth}_{\rho}(\rho(j))}X_{n,j,k}\leq t\log n\right)$
$\hskip 60.0pt=\lim_{n\to\infty}\sum_{k=1}^{m-1}\mathbb{P}\left(\sum_{i=1}^{k}X_{n,j,k}\leq t\log n\mid\operatorname{depth}_{\rho}(\rho(j))=k\right)\mathbb{P}(\operatorname{depth}_{\rho}(\rho(j))=k)$
$\hskip 60.0pt=\sum_{k=1}^{m-1}\lim_{n\to\infty}\mathbb{P}\left(\sum_{i=1}^{k}X_{n,j,k}\leq t\log n\right)\mathbb{P}(\operatorname{depth}_{\rho}(\rho(j))=k)$
$\hskip 60.0pt=\sum_{k=1}^{m-1}\mathbb{P}(k\leq t)\mathbb{P}(\operatorname{depth}_{\rho}(\rho(j))=k)$
$\hskip 60.0pt=\sum_{k=1}^{m-1}\mathbb{P}(\operatorname{depth}_{\rho}(\rho(j))\leq t,\operatorname{depth}_{\rho}(\rho(j))=k)$
$\hskip 60.0pt=\mathbb{P}(\operatorname{depth}_{\rho}(\rho(j))\leq t).$
Building on top of these pieces and now also using the Continuous mapping theorem (applied to the function ${\mathbf{z}}\mapsto\max_{j}z_{j}$) and Slutsky’s theorem (Theorem 5), it follows then $h(\mathcal{T}(\pi))/\log n=\max_{j}z_{j}/\log n$ converges in distribution to
$c^{*}+\max_{j}\operatorname{depth}_{\rho}(\rho(j))=c^{*}+h(\mathcal{T}(\rho)).$
Combining all of these pieces and taking expectations with respect to $\rho$ to undo the conditioning on $\rho$ then completes the proof of Theorem 2, noting also that $\log nm=\log n+\log m=\log n\cdot(1+o(1))$ when $m$ is fixed. ∎
A first takeaway from Theorem 2 is that one wreath product or Kronecker product is sufficient to increase the height of a block BST from the uniform case. Thus, Theorem 2 is a generalization of Theorem 1, which is now included as a sub-case when $m=1$ (note $h_{1}=0$). For $m=2$ then $h(\mathcal{T}(\rho))=1$ (deterministically), so that $h_{m,n}:=h(\mathcal{T}(\pi))$ limits to $c^{*}+1$ when scaled by $\log n$. For $m\geq 3$, then $h_{m}\geq 1$ is now random (e.g., $h_{3}\sim 1+\operatorname{Bern}(\frac{2}{3})$), so that $c^{*}+h_{m}$ is a random distributional limit for $h_{m,n}/\log n$ for $m\geq 3$ that is thus strictly larger than $c^{*}$. Note also though the BST height remains proportional to the logarithmic length of the input vector (i.e., $\log nm$).
This presents several natural directions for further study:
  • What is the asymptotic behavior of $h_{m,n}$ if $m$ grows? What happens if $m$ grows while $n$ is fixed? If $m=\Omega(n)$?
  • How much can the height of a block BST grow when the input permutation is generated through repeated applications of wreath or Kronecker products? Does this remain $O(\log nm)$?
The remainder of this paper, starting in Section 3, will focus on addressing the second set of questions. The main set of block BST generating permutations of length $N=2^{n}$ we will consider are initialized by $S_{2}$ and then formed either by using iterated wreath products or Kronecker products. This is the smallest set of (non-trivial) permutations generated using only wreath or Kronecker products. These align then exactly with the nonsimple and simple butterfly permutation groups, $\operatorname{B}_{n},\operatorname{B}_{n,s}\subset S_{N}$ (see Sections 3.1 and 3.2).
For the remainder of the current section, we will partially address the first set of questions. Theorem 2 fully answers the questions regarding the asymptotic behavior of $h_{m,n}$ when $m$ is fixed and $n$ grows, so we will next consider the case when $n$ is fixed and $m$ grows. Again, this generalizes Theorem 2 in the uniform setting, as this can now be subsumed as the case when $n=1$. For the general case, we will first establish a lower boundary:
Proposition 1.
Let $\pi\sim\operatorname{Unif}(S_{n}\wr S_{m})$ or $\pi\sim\operatorname{Unif}(S_{m}\otimes S_{n})$, and $h_{n,m}=h(\mathcal{T}(\pi))$.Then
$\lim_{m\to\infty}\mathbb{P}(h_{n,m}\geq c^{*}H_{n}^{(1)}\log m)=1.$
Proof.
This is equivalent to finding a specific path in $\pi$ of depth at least $c^{*}H_{n}^{(1)}$ asymptotically. Let $\rho\in S_{m}$ denote the induced block permutation from $\pi$. We will construct such a path by focusing first on an $m$-block of maximal depth in $\rho$, and then choosing a node of maximal depth within this block. Let $\rho(j)$ denote a specific node in $\mathcal{T}(\rho)$ of maximal depth, i.e., $\operatorname{depth}_{\rho}(\rho(j))=h(T(\rho))=:h_{m}$. Let $v$ denote a node within $\rho(j)$ of maximal depth in $\mathcal{T}(\pi_{\rho(j)})$. By Corollary 2 and following the notation from the proof of Theorem 2, it follows then
$Z_{n,m}:=\operatorname{depth}_{\pi}(v)=h_{n}+\sum_{k=1}^{h_{m}}X_{n,j,k}$
where $X_{n,j,k}\sim\Upsilon_{n}$ iid for each $k$ when $n,j$ is fixed, and we write $h_{n}:=h(\mathcal{T}(\pi_{\rho(j)})$. In particular, we note $h_{n,m}\geq Z_{n,m}$ trivially, while also by the Subgroup algorithm $\rho\sim\operatorname{Unif}(S_{m})$ and $\pi_{j}^{\prime}\sim\operatorname{Unif}(S_{n})$. To establish Proposition 1, it now suffices to show $Z_{n,m}/\log m\to_{p}c^{*}H_{n}^{(1)}$ as $m\to\infty$.
For $n$ fixed, let
$W_{m}:=\frac{1}{h_{m}}\sum_{k=1}^{h_{m}}X_{n,j,k}=\overline{(\Upsilon_{n})}_{h_{m}},$
i.e., a sample mean with $h_{m}$ iid samples of $\Upsilon_{n}$ (where we again emphasize that $h_{m}$ is random). Then
$Z_{n,m}=h_{n}+h_{m}W_{m}.$
If $m\to\infty$, then $h_{n}/\log m\to_{p}0$, $h_{m}/\log m\to_{p}c^{*}$, and now we claim $W_{m}\to_{p}\mathbb{E}\Upsilon_{n}=H_{n}^{(1)}$. Using also Slutsky’s theorem as well as the equivalence of convergence in probability and distribution when the limit is constant, this would then yield $Z_{n,m}/\log m\to_{p}c^{*}H_{n}^{(1)}$ when $n$ is fixed, as desired. Now $W_{m}\to_{p}H_{n}^{(1)}$ follows immediately from the Strong Law of Large Numbers (SLLN) applied to the random sample mean $W_{m}$ along with the fact $h_{m}=\omega(1)$ (since $h_{m}\geq\lfloor\log_{2}m\rfloor$). ∎
We conjecture additionally that $h_{n,m}/\log m\to_{p}c^{*}H_{n}^{(1)}$ as well when $n$ is fixed and $m$ grows. To establish then a matching upper bound with the same limit would require a more involved update of Devroye’s original corresponding argument that connected heights in BSTs to depths within a fully balanced tree, and established strong results from extreme values in branching processes. Hence, noting also $H_{n}^{(1)}/\log n\to 1$ as $n\to\infty$, then this suggests that $h_{n,m}$ should limit to $c^{*}$ when both $m,n$ grow, but now scaling by $\log n\log m$:
Conjecture 1.
If $\pi\sim\operatorname{Unif}(S_{n}\wr S_{m})$ or $\pi\sim\operatorname{Unif}(S_{m}\otimes S_{n})$, then $h_{n,m}/\log n\log m\to_{p}c^{*}$ as $\min(n,m)\to\infty$.
In the least, Theorem 2 and Proposition 1 together yield a matching lower bound for 1.
3 Heights of butterfly trees
We will now address the question of the height of a random BST generated using a permutation of length $N=2^{n}$ formed using iterated wreath and Kronecker products, initiating with $S_{2}$. As seen in Theorem 2, a single use of a wreath or Kronecker product for the generating permutation suffices to increase the asymptotic height of the associated BST, from $c^{*}\approx 4.311$ when scaled by $\log n$ in the uniform $S_{n}$ case to now $c^{*}+h_{m}$ when scaled by $\log n$ still in the uniform $S_{n}\wr S_{m}$ or $S_{m}\otimes S_{n}$ case, when $m$ is fixed. We are now interested in exploring how much this height can increase when using iterated products of permutations, focusing on the simplest recursive cases that initiate with $S_{2}$ and so generate permutations of length $N=2^{n}$. These permutations then comprise precisely the simple butterfly permutations, $\operatorname{B}_{n,s}=S_{2}^{\otimes n}$, and the nonsimple butterfly permutations, $\operatorname{B}_{n}=S_{2}^{\wr n}$ (see Section 1.3 for associated definitions). A butterfly tree is then the BST $\mathcal{T}(\boldsymbol{\pi}_{n})$ when $\boldsymbol{\pi}_{n}\in S_{N}$ is a butterfly permutation. Write $h_{n}^{\operatorname{B}}=h(\mathcal{T}(\boldsymbol{\pi}_{n}))$ then for the height of a butterfly tree, and similarly write $\ell_{n}^{\operatorname{B}}$ and $r_{n}^{\operatorname{B}}$ for the lengths of the top-left and top-right edges of $\mathcal{T}(\boldsymbol{\pi}_{n})$, which we further recall are associated to the left-to-right minima and maxima paths of $\boldsymbol{\pi}_{n}$.
The butterfly permutations have been studied extensively in prior works. As previously mentioned, $\operatorname{B}_{n}$ has been of particular interest under the guise of its comprising a 2-Sylow subgroup of $S_{2^{n}}$ (see [1], with subsequent applications across group-theoretic and probabilistic settings). In preceding work [36], we additionally focused on the questions of the number of cycles and the LIS for butterfly permutations, written respectively as $C(\pi)$ and $\operatorname{LIS}(\pi)$ for permutation $\pi$. For simple butterfly permutations, explicit recurrence relations yield precise distributional results:
Theorem 7 ([36]).
If $\boldsymbol{\pi}_{n}\sim\operatorname{Unif}(\operatorname{B}_{n,s})$, with $\boldsymbol{\pi}_{n}=\bigotimes_{j=1}^{n}\pi_{j}$ for $\pi_{j}\in S_{2}$, then $\frac{2}{N}C(\boldsymbol{\pi}_{n})-1=\prod_{j=1}^{n}\mathds{1}(\pi_{j}=12)\sim\operatorname{Bern}(\frac{1}{N})$ while $\log_{2}\operatorname{LIS}(\boldsymbol{\pi}_{n})=\sum_{j=1}^{n}\mathds{1}(\pi_{j}=12)\sim\operatorname{Binom}(n,\frac{1}{2})$ and $\operatorname{LIS}(\boldsymbol{\pi}_{n})\cdot\operatorname{LDS}(\boldsymbol{\pi}_{n})=N$, with then $\frac{2}{N}C(\boldsymbol{\pi}_{n})\to_{p}1$ and $(\log_{2}\operatorname{LIS}(\boldsymbol{\pi}_{n})-n/2)/\sqrt{n/4}\to_{d}N(0,1)$ as $n\to\infty$.
For nonsimple butterfly permutations, similar but more intricate recursions arise. The LIS obeys a branching-type recurrence, while the cycle count admits a full (non-asymptotic and asymptotic) distributional description:
Theorem 8 ([36]).
If $\boldsymbol{\pi}_{n}\sim\operatorname{Unif}(\operatorname{B}_{n})$, then $\operatorname{LIS}(\boldsymbol{\pi}_{n})\sim X_{n}$ recursively defined where $X_{1}=1$ and $X_{n+1}\sim(X_{n}+X_{n}^{\prime})\eta_{n}+\max(X_{n},X_{n}^{\prime})(1-\eta_{n})$ for iid $X_{n}\sim X_{n}^{\prime}$ independent of $\eta_{n}\sim\operatorname{Bern}(\frac{1}{2})$ for $n\geq 2$, and there exist constants $\alpha=\log_{2}(3/2)\approx 0.58496$ and $\beta\approx 0.83255$ such that $N^{\alpha}\leq\mathbb{E}\operatorname{LIS}(\boldsymbol{\pi}_{n})\leq N^{\beta}$. Moreover, $C(\boldsymbol{\pi}_{n})\sim Y_{n}$ recursively defined where $Y_{1}=1$ and $Y_{n+1}\sim Y_{n}+\eta_{n}Y_{n}^{\prime}$ for iid $Y_{n}\sim Y_{n}^{\prime}$ independent of $\eta_{n}\sim\operatorname{Bern}(\frac{1}{2})$, with $C(\boldsymbol{\pi}_{n})/N^{\alpha}\to_{d}W$ where $W$ has positive support that is uniquely determined by its positive integer moments $m_{k}=\mathbb{E}W^{k}$ satisfying $m_{0}=m_{1}=1$ and for $k\geq 2$ then
$m_{k}=\frac{\lambda-1}{\lambda^{k}-1}\sum_{j=1}^{k-1}\binom{k}{j}m_{j}m_{k-j}$
for $\lambda=2^{\alpha}=3/2$
Remark 1.
As shown in Theorem 8, both the LIS and number of cycles for permutations generated via iterated wreath products exhibit significantly larger average values than under uniform permutations. Specifically, the average LIS and number of cycles increase from orders $N^{1/2}$ and $\log N$, respectively, to at least $N^{\alpha}\approx N^{0.58496}$. For the number of cycles, this shift from logarithmic to polynomial growth precisely mirrors the corresponding increase in height for BSTs built from iterated wreath products versus those built from uniform permutations (see Theorem 4).
In this work, we adapt and extend the analytical techniques from [36] to study the height $h_{n}^{\operatorname{B}}$ of butterfly trees. We establish key connections between the left and right edge lengths ($r_{n}^{\operatorname{B}},\ell_{n}^{\operatorname{B}}$) and the classical permutation statistics $C(\pi)$ and $\operatorname{LIS}(\pi)$ (the main focus of [36]), which form the core of our approach to deriving distributional properties of butterfly tree heights. By further leveraging the recursive structure of butterfly trees, we derive exact distributional results for the height in the simple case and establish strong asymptotic bounds in the nonsimple case, highlighting how these recursive and combinatorial elements shape the growth of butterfly tree heights.
Of particular interest is the contrast with the uniform case (Section 1.2), where the expected BST height grows logarithmically. As shown in Theorem 2, this logarithmic behavior persists even after applying a single wreath or Kronecker product. However, for butterfly permutations formed by iterated products, the expected BST height grows polynomially, with $\mathbb{E}h_{n}^{\operatorname{B}}=\Omega(N^{\alpha})$ (see Theorems 3 and 4). This represents a significant departure from the classical setting and underscores the combinatorial complexity inherent in butterfly permutations within the BST framework.
3.1 Simple butterfly trees: $\mathcal{T}^{\operatorname{B}}_{n,s}$
We will first consider simple butterfly trees, $\mathcal{T}^{\operatorname{B}}_{n,s}$, generated using uniform simple butterfly permutations $\boldsymbol{\pi}_{n}=\bigotimes_{j=1}^{n}\pi_{j}\sim\operatorname{Unif}(\operatorname{B}_{n,s})$. Figure 4 shows two instances of simple butterfly trees, since we further note $2143=12\otimes 21$. Hence, $21436587=12\otimes 12\otimes 21$ and $65872143=21\otimes 12\otimes 21$ are each simple butterfly permutations from $\operatorname{B}_{3,s}$. When building up $\operatorname{B}_{n,s}$ using Kronecker products of either 12 or 21, then we can write $\boldsymbol{\pi}_{n+1}=\pi_{n+1}\otimes\boldsymbol{\pi}_{n}\in\operatorname{B}_{n+1,s}$ recursively as a Kronecker product of $\pi_{n+1}\in S_{2}=\operatorname{B}_{1,s}$ and $\boldsymbol{\pi}_{n}\in\operatorname{B}_{n+1,s}$ (cf. Figure 4).
Similar to the LIS question for simple butterfly permutations, Theorem 3 shows we again are able to attain exact nonasymptotic and asymptotic distributional descriptions for the heights of simple butterfly trees, $h_{n}^{\operatorname{B}}$. The proof of Theorem 3 follows from a series of recursive relationships relating $\boldsymbol{\pi}_{n+1}=\pi_{n+1}\otimes\boldsymbol{\pi}_{n}$ directly to properties each of $\pi_{n+1}$ and $\boldsymbol{\pi}_{n}$, which we will establish through a series of lemmas.
We begin by highlighting a structural property of simple butterfly trees: their height decomposes exactly into the sum of the lengths of their top-left and top-right edge paths. This clean decomposition reflects a recursive geometry that will later connect to the LIS and LDS (longest decreasing subsequence) of $\boldsymbol{\pi}_{n}$. Namely, we have:
Lemma 2.
Let $\boldsymbol{\pi}_{n}\in\operatorname{B}_{n,s}$. Then $h(\mathcal{T}(\boldsymbol{\pi}_{n}))=\ell(\mathcal{T}(\boldsymbol{\pi}_{n}))+r(\mathcal{T}(\boldsymbol{\pi}_{n}))$.
Proof.
For the sake of compactness, write $h_{n}^{\operatorname{B}}=h(\mathcal{T}(\boldsymbol{\pi}_{n}))$, $\ell_{n}^{\operatorname{B}}=\ell(\mathcal{T}(\boldsymbol{\pi}_{n}))$, and $r_{n}^{\operatorname{B}}=r(\mathcal{T}(\boldsymbol{\pi}_{n}))$. We will establish this by induction on $n$ for $\boldsymbol{\pi}_{n}\in\operatorname{B}_{n,s}$. The $n=1$ case is immediate, as $\ell_{1}^{\operatorname{B}}=\mathds{1}(\pi_{1}=12)$ and $r_{1}^{\operatorname{B}}=\mathds{1}(\pi_{1}=21)=1-\ell_{1}^{\operatorname{B}}$ as each can be either of 0 or 1 depending on the value $\pi_{1}\in S_{2}$ (see Figure 3(a-b) for instances of $\mathcal{T}(12)$ and $\mathcal{T}(21)$). Hence, noting the height is deterministically equal to 1 in either case, we then have $h_{1}^{\operatorname{B}}=1=\ell_{1}^{\operatorname{B}}+r_{1}^{\operatorname{B}}$.
Now assume the result holds for $\operatorname{B}_{m,s}$ for $m\leq n$, and let $\boldsymbol{\pi}_{n+1}\in\operatorname{B}_{n+1,s}$. Let $\pi_{n+1}\in S_{2}$ and $\boldsymbol{\pi}_{n}\in\operatorname{B}_{n,s}$ such that $\boldsymbol{\pi}_{n+1}=\pi_{n+1}\otimes\boldsymbol{\pi}_{n}$. First, consider the case when $\pi_{n+1}=12$, then $\mathcal{T}(\boldsymbol{\pi}_{n+1})$ is constructed by gluing together two copies of $\mathcal{T}(\boldsymbol{\pi}_{n})$ connecting from the bottom of the top-right edge of the top copy of $\mathcal{T}(\boldsymbol{\pi}_{n})$ to the root of the right-child copy of $\mathcal{T}(\boldsymbol{\pi}_{n})$ (i.e., the sub-tree $\mathcal{T}(\boldsymbol{\pi}_{n}+N)$). Hence, the top-right edge of $\mathcal{T}(\boldsymbol{\pi}_{n+1})$ is comprised of the top-right edges of the two copies of $\mathcal{T}(\boldsymbol{\pi}_{n})$ along with the extra edge gluing together the two sub-trees (i.e., $r_{n+1}^{\operatorname{B}}=2\cdot r_{n}^{\operatorname{B}}+1$), while the top-left edge is comprised of only the top-left edge of the parent copy of $\mathcal{T}(\boldsymbol{\pi}_{n})$ (i.e., $\ell_{n+1}^{\operatorname{B}}=\ell_{n}^{\operatorname{B}}$). In particular, then we note the height of $\mathcal{T}(\boldsymbol{\pi}_{n+1})$ is necessarily determined by the right-sub-tree of the root, and is thus determined by the path traveling the top-right edge of the parent copy, passing over the gluing edge, followed by a maximal path in the lower child (i.e., $h_{n+1}^{\operatorname{B}}=r_{n}^{\operatorname{B}}+1+h_{n}^{\operatorname{B}}$).
Similarly, when $\pi_{n+1}=21$, then the top-left edge of $\mathcal{T}(\boldsymbol{\pi}_{n+1})$ is comprised of the top-left edges of the sub-trees of copies of $\mathcal{T}(\boldsymbol{\pi}_{n})$ along with the extra edge gluing together the parent and now left-child block sub-tree while the top-right edge is comprised of only the top-right edge of the parent copy of $\mathcal{T}(\boldsymbol{\pi}_{n})$, with then the height of the entire tree determined by the top-right edge length of the parent block tree added to the height of the child block tree. To summarize, we have
$\ell_{n+1}^{\operatorname{B}}$ =$=\left\{\begin{array}[]{ll}\ell_{n}^{\operatorname{B}},&\pi_{n+1}=12\\ 2\cdot\ell_{n}^{\operatorname{B}}+1,&\pi_{n+1}=21\end{array}\right.$
$r_{n+1}^{\operatorname{B}}$ =$=\left\{\begin{array}[]{ll}2\cdot r_{n}^{\operatorname{B}}+1,&\pi_{n+1}=12\\ r_{n}^{\operatorname{B}},&\pi_{n+1}=21\\ \end{array}\right.$
$h_{n+1}^{\operatorname{B}}$ =$=\left\{\begin{array}[]{ll}r_{n}^{\operatorname{B}}+1+h_{n}^{\operatorname{B}},&\pi_{n+1}=12\\ \ell_{n}^{\operatorname{B}}+1+h_{n}^{\operatorname{B}},&\pi_{n+1}=21\\ \end{array}\right.$
This can be condensed into defining the recursion on the vector $(h_{n}^{\operatorname{B}},\ell^{\operatorname{B}},r_{n}^{\operatorname{B}})$ as
$(h_{n+1}^{\operatorname{B}},\ell_{n+1}^{\operatorname{B}},r_{n+1}^{\operatorname{B}})$ $=(h_{n}^{\operatorname{B}},\ell_{n}^{\operatorname{B}},r_{n}^{\operatorname{B}})+\left\{\begin{array}[]{ll}(r_{n}^{\operatorname{B}}+1)(1,0,1),&\pi_{n+1}=12\\ (\ell_{n}^{\operatorname{B}}+1)(1,1,0),&\pi_{n+1}=21\end{array}\right.$ (6)
Using this recursive structure on $(h_{n}^{\operatorname{B}},\ell_{n}^{\operatorname{B}},r_{n}^{\operatorname{B}})$, it follows inductively that
$(h_{n}^{\operatorname{B}},\ell_{n}^{\operatorname{B}},r_{n}^{\operatorname{B}})\in\operatorname{Span}\{(1,0,1),(1,1,0)\}.$
An immediate consequence from this is necessarily
$h_{n}^{\operatorname{B}}=\ell_{n}^{\operatorname{B}}+r_{n}^{\operatorname{B}},$
as $x=y+z$ holds for $(x,y,z)\in\operatorname{Span}\{(1,0,1),(1,1,0)\}$ since this obviously holds for each $(1,0,1)$ and $(1,1,0)$. ∎
Remark 2.
Lemma 2 is in stark contrast to the uniform permutation case $\pi\sim\operatorname{Unif}(S_{n})$, where each height and left-to-right maxima/minima scale as $\Theta(\log n)$ with differently limiting coefficients, so that $h_{n}-(\ell_{n}+r_{n})\gtrsim(c^{*}-2)\log n+o(\log n)=\omega(1)$ typically.
Now note also from Theorem 7, since $N=\operatorname{LIS}(\boldsymbol{\pi}_{n})\cdot\operatorname{LDS}(\boldsymbol{\pi}_{n})$ while $\log_{2}\operatorname{LIS}(\boldsymbol{\pi}_{n})=\sum_{j=1}^{n}\mathds{1}(\pi_{j}=12)\sim\operatorname{Binom}(n,1/2)$, then also $\log_{2}\operatorname{LDS}(\boldsymbol{\pi}_{n})=\sum_{j=1}^{n}\mathds{1}(\pi_{j}\neq 12)=n-\log_{2}\operatorname{LIS}(\boldsymbol{\pi}_{n})$. (This can be proved directly, as seen in [36], by considering the LIS as a down-right path connecting 1’s in the associated permutation matrix for $\boldsymbol{\pi}_{n}$, and so analogously the LDS as an up-right path connecting 1’s; inductively using straightforward recursion on the block matrix permutation form, one of each of the LIS and LDS are either maintained or doubled for each intermediate Kronecker product step.) It suffices to observe that this same recursion is satisfied by $\ell(\mathcal{T}(\boldsymbol{\pi}_{n}))$ and $r(\mathcal{T}(\boldsymbol{\pi}_{n}))$, which the above inductive construction confirms. It then follows immediately:
Lemma 3.
Let $\boldsymbol{\pi}_{n}\in\operatorname{B}_{n,s}$. Then $\operatorname{LIS}(\boldsymbol{\pi}_{n})=\ell(\mathcal{T}(\boldsymbol{\pi}_{n}))+1$ and $\operatorname{LDS}(\boldsymbol{\pi}_{n})=r(\mathcal{T}(\boldsymbol{\pi}_{n}))+1$.
Hence, together this yields the first statement from Theorem 3:
$h_{n}^{\operatorname{B}}\sim 2^{X_{n}}+2^{n-X_{n}}-2,\qquad X_{n}\sim\operatorname{Binom}(n,1/2)$ (7)
Since $\mathbb{E}2^{X_{n}}=2^{-n}\sum_{k=0}^{n}2^{k}\binom{n}{k}=(3/2)^{n}$ and $X_{n}\sim n-X_{n}$, it follows immediately
$\mathbb{E}h_{n}^{\operatorname{B}}=2\left(\frac{3}{2}\right)^{n}-2=2\cdot N^{\alpha}\cdot(1+o(1)),\qquad\alpha=\log_{2}(3/2)\approx 0.58496.$
[…]
$k$ $f(k)=2^{k}+2^{10-k}-2$$\#\{\boldsymbol{\pi}\in\operatorname{B}_{10,s}:h(\mathcal{T}(\boldsymbol{\pi}))=f(k)\}$
0 1023 2
1 512 20
2 258 90
3 134 240
4 78 420
5 62 252
[…]
Remark 3.
Recall in the uniform permutation case $\pi_{n}\sim\operatorname{Unif}(S_{n})$, then $h(\mathcal{T}(\pi_{n}))/\log n\to_{p}c^{*}\approx 4.311$ (the unique solution to $x\log(2e/x)=1$ for $x\geq 2$) while $\operatorname{LIS}(\pi_{n})/\sqrt{n}\to_{p}2$. In particular, $h(\mathcal{T}(\pi_{n}))$ is asymptotically dominated by $\operatorname{LIS}(\pi_{n})$. For the simple binary butterfly permutation case, this is a different story, since
$\mathbb{E}h(\mathcal{T}(\pi_{n}))=\mathbb{E}\operatorname{LIS}(\pi_{n})+\mathbb{E}\operatorname{LDS}(\pi_{n})-2=2\mathbb{E}\operatorname{LIS}(\pi_{n})-2$
using $\operatorname{LIS}(\pi_{n})\sim\operatorname{LDS}(\pi_{n})$, so that $\mathbb{E}h(\mathcal{T}(\pi_{n}))=\Theta(\mathbb{E}\operatorname{LIS}(\pi_{n}))$.
Remark 4.
We note the distribution given in (7) matches exactly the degree distribution of a uniformly sampled vertex from the comparability graph of the Boolean lattice with $n$-element ground set. Figure 5 displays the adjacency matrix for this graph. Remarkably, its sparsity pattern coincides with that of $L+U$, where $PB=LU$ is the the GEPP factorization of a simple butterfly matrix $B$—up to the diagonal entries, which are zero in the adjacency matrix and nonzero in $L+U$. (Interestingly, the visualization appears to manifest several butterfly-like forms.) In this factorization, the permutation matrix factor $P$ corresponds to a simple butterfly permutation. Future work can explore additional connections between random graphs and GEPP.
Figure 5: Sparsity pattern map for the adjacency matrix for the comparability graph of the Boolean lattice of size $2^{12}$
The remaining part of Theorem 3 requires establishing a distributional limit for $(\log_{2}h(\mathcal{T}(\pi_{n})-n/2)/(\sqrt{n}/2)\to_{d}|Z|$ as $n\to\infty$ to $|Z|$ for $Z\sim N(0,1)$ for uniform simple butterfly permutations $\pi_{n}$. This follows using standard results from probability theory:
Let $Z_{n}=2(X_{n}-n/2)/\sqrt{n}$ with $Z_{n}\to_{d}Z\sim N(0,1)$ by the Central Limit Theorem (CLT), with $h_{n}^{\operatorname{B}},X_{n}$ defined as in (7). Let $R_{n}=h_{n}^{\operatorname{B}}+2$. Then
$W_{n}:=2^{-n/2}R_{n}=2^{X_{n}-n/2}+2^{n/2-X_{n}}=2\cosh(\ln 2\sqrt{n}|Z_{n}|/2)=2^{\sqrt{n}|Z_{n}|/2}(1+2^{-\sqrt{n}|Z_{n}|}),$
and now let
$Q_{n}=W_{n}^{2/\sqrt{n}}:=A_{n}B_{n}^{2/\sqrt{n}}$
where $A_{n}=2^{|Z_{n}|}$ and $B_{n}=1+2^{-\sqrt{n}|Z_{n}|}$. Note first $B_{n}\to_{p}1$. This is straight forward, but to include the details: Fix $\epsilon>0$, and note
$\mathbb{P}(|B_{n}-1|=2^{-\sqrt{n}|Z_{n}|}>\epsilon)=\mathbb{P}\left(|Z_{n}|<\frac{\log_{2}\frac{1}{\epsilon}}{\sqrt{n}}\right)$
If $\epsilon\geq 1$, then this evaluates directly to 0 since $|Z_{n}|\geq 0$ while $\log_{2}\frac{1}{\epsilon}\leq 0$, so we only need to consider $\epsilon\in(0,1)$. Fix $\delta>0$ and let $N_{\delta}$ be such that $\frac{\log_{2}\frac{1}{\epsilon}}{\sqrt{n}}<\delta$ if $n\geq N_{\delta}$. It follows then
$\mathbb{P}(|B_{n}-1|>\epsilon)\leq\mathbb{P}(|Z_{n}|<\delta)$
for $n\geq N_{\delta}$, so that
$\limsup_{n\to\infty}\mathbb{P}(|B_{n}-1|>\epsilon)\leq\mathbb{P}(|Z|<\delta)=2\Phi(\delta)-1=o(\delta)$
by the CLT, with $\Phi$ denoting the cumulative distribution function for $Z$; it follows $B_{n}\to_{p}1$ by taking $\delta\to 0$. In particular, then also $B_{n}^{2/\sqrt{n}}\to_{p}1$ by other standard probabilistic methods. Since $A_{n}\to_{d}2^{|Z|}$ (by the Continuous mapping theorem), it follows then by Slutsky’s theorem that
$Q_{n}=A_{n}B_{n}^{2/\sqrt{n}}\to_{d}2^{|Z|},$
where we recall
$Q_{n}=W_{n}^{2/\sqrt{n}}=(2^{-n/2}R_{n})^{2/\sqrt{n}}=[2^{-n/2}(h_{n}^{\operatorname{B}}+2)]^{2/\sqrt{n}}.$
Hence, also (by the Continuous mapping theorem)
$\log_{2}Q_{n}=2(\log_{2}(h_{n}^{\operatorname{B}}+2)-n/2)/\sqrt{n}\to_{d}|Z|.$
Replacing $h_{n}^{\operatorname{B}}+2$ directly with just $h_{n}^{\operatorname{B}}$ above then has the same distributional limit (via Slutsky’s theorem), i.e.,
$2(\log_{2}h_{n}^{\operatorname{B}}-n/2)/\sqrt{n}\to_{d}|Z|.$
(This mirrors the same normalization that determined $Z_{n}$ from $X_{n}$ for the CLT application above.) Equivalently, this shows
$[2^{-n/2}h_{n}^{\operatorname{B}}]^{2/\sqrt{n}}=[h_{n}^{\operatorname{B}}/\sqrt{N}]^{2/\sqrt{n}}=[(h_{n}^{\operatorname{B}})^{2}/N]^{1/\sqrt{n}}\to_{d}2^{|Z|}.$
Combining all of these results now completes the proof of Theorem 3.
3.2 Nonsimple butterfly trees: $\mathcal{T}^{\operatorname{B}}_{n}$
Now we turn to nonsimple butterfly trees, $\mathcal{T}_{n}^{\operatorname{B}}$, which are generated using $\boldsymbol{\pi}\sim\operatorname{Unif}(\operatorname{B}_{n})$. Unlike the simple case, where the generating simple permutations are constructed using Kronecker products, the nonsimple case involves wreath product structures. For instance, instead of structures like $12\otimes\pi_{1}=\pi_{1}\oplus\pi_{1}$ or $21\otimes\pi_{1}=\pi_{1}\ominus\pi_{1}$, the nonsimple setting allows for different subblocks, such as $\pi_{1}\oplus\pi_{2}$ and $\pi_{1}\ominus\pi_{2}$, with $\pi_{1},\pi_{2}$ not necessarily equal. These can be alternatively written as: $\pi_{1}\oplus\pi_{2}=(12\otimes\pi_{1})(1\oplus\pi_{1}^{-1}\pi_{2})$. This greater structural flexibility complicates the analysis of the height $h_{n}^{\operatorname{B}}=h(\mathcal{T}(\boldsymbol{\pi}))$ for $\boldsymbol{\pi}\sim\operatorname{Unif}(\operatorname{B}_{n})$, making a full distributional characterization (cf. Theorem 3) less tractable. Instead, we focus on bounding the expected height. We establish power-law bounds so that $cN^{\alpha}\cdot(1+o(1))\leq E[h_{n}^{B}]\leq dN^{\beta}\cdot(1+o(1))$ for $\beta\approx 0.913189$, as detailed in Theorem 4.
For a permutation of these block structures, these have corresponding BST $\mathcal{T}(\pi)$ that can similarly be constructed recursively by replacing each node of an external BST with an internal BST at each generating step.
$\mathcal{T}(\pi_{1})$$\mathcal{T}(\pi_{2})$
(a) $\mathcal{T}(\pi_{1}\oplus\pi_{2})$
$\mathcal{T}(\pi_{1})$$\mathcal{T}(\pi_{2})$
(b) $\mathcal{T}(\pi_{1}\ominus\pi_{2})$
65827143910111213141516
(c) $\mathcal{T}(9,10,11,12,13,14,15,16,6,5,8,7,2,1,4,3)=T(65872143\ominus 12345678)$
Figure 6: $\mathcal{T}(\boldsymbol{\pi})$ for $\boldsymbol{\pi}=(\rho\otimes\operatorname{id})(\pi_{1}\oplus\pi_{2})$ for $\rho=12$ (in (a)) and $\rho=21$ (in (b)); while (c) shows a particular instance of the form $\mathcal{T}(\pi_{1}\ominus\pi_{2})$ where $\pi_{1}\neq\pi_{2}$
To analyze heights for these corresponding butterfly trees, we will leverage the recursive construction and relationships between each corresponding level of permutations. Let $\boldsymbol{\pi}\in B_{n+1}$, and so let $\rho\in S_{2}$ and $\pi_{1},\pi_{2}\in B_{n}$ such that $\boldsymbol{\pi}=(\rho\otimes\operatorname{id})(\pi_{1}\oplus\pi_{2})$. Let $H_{n+1}=h(\mathcal{T}(\boldsymbol{\pi}))$, $H_{n}=h(\mathcal{T}(\pi_{1}))$, $H_{n}^{\prime}=h(\mathcal{T}(\pi_{2}))$, $L_{n+1}=\ell(\mathcal{T}(\boldsymbol{\pi}))$, $L_{n}=\ell(\mathcal{T}(\pi_{1}))$, $L_{n}^{\prime}=\ell(\mathcal{T}(\pi_{2}))$, $R_{n+1}=r(\mathcal{T}(\boldsymbol{\pi}))$, $R_{n}=r(\mathcal{T}(\pi_{1}))$, and $R_{n}^{\prime}=r(\mathcal{T}(\pi_{2}))$. Figure 6 shows the two structures for $\mathcal{T}(\boldsymbol{\pi})$ depending on the values of $\rho\in S_{2}$. The corresponding values for $\boldsymbol{\pi}$ depend on $\rho\in S_{2}$ along with the corresponding values for $\pi_{1}$ and $\pi_{2}$. For example, for $\boldsymbol{\pi}=\pi_{1}\ominus\pi_{2}$ (as in Figure 6b, where $\rho=21$) then $R_{n+1}=R_{n}^{\prime}$ and $L_{n+1}=L_{n}^{\prime}+1+L_{n}$, while $H_{n+1}=\max(L_{n}^{\prime}+1+H_{n},H_{n}^{\prime})$. Figure 6 shows an instance where $H_{n+1}=H_{n}^{\prime}=7>L_{n}^{\prime}+1+H_{n}=0+1+4=5$; note for further comparison, the included $\boldsymbol{\pi}=\pi_{1}\ominus\pi_{2}$ has $\operatorname{LIS}(\pi)=8$ while $\operatorname{LDS}(\pi)=5$. In general, then $(H_{n},L_{n},R_{n})$ satisfy the following recursion formulae:
$L_{n+1}$ =$=\left\{\begin{array}[]{ll}L_{n},&\rho=12\\ L_{n}^{\prime}+1+L_{n},&\rho=21\end{array}\right.$ (10)
$R_{n+1}$ =$=\left\{\begin{array}[]{ll}R_{n}+1+R_{n}^{\prime},&\rho=12\\ R_{n}^{\prime},&\rho=21\end{array}\right.$ (13)
$H_{n+1}$ =$=\left\{\begin{array}[]{ll}\max(R_{n}+1+H_{n}^{\prime},H_{n}),&\rho=12\\ \max(L_{n}^{\prime}+1+H_{n},H_{n}^{\prime}),&\rho=21\end{array}\right.$ (16)
From (16), it’s clear $(H_{n},L_{n},R_{n})$ cannot satisfy a simple recursion as seen for simple binary butterfly trees (cf. (6)).
The appearance of the $\max$ in the heights is the main obstacle in computing the first moment. However, the recursive formulae for each $L_{n}$ and $R_{n}$ allow direct analysis, which also compares directly to the number of cycles of a nonsimple (binary) butterfly permutation, $C(\pi_{n})\sim Y_{n}$ where $Y_{0}=1$ and $Y_{n+1}\sim Y_{n}+\eta_{n}Y_{n}^{\prime}$ for $Y_{n}\sim Y_{n}^{\prime}$ iid, independent of $\eta_{n}=\mathds{1}(\rho=12)\sim\operatorname{Bern}(1/2)$, as shown previously in Theorem 8. Now (10) and (13) yield $L_{n}$ and $R_{n}$ satisfy the same recurrence (up to shifting). Moreover, noting $C(\boldsymbol{\pi})=C(\boldsymbol{\pi_{2}})+\mathds{1}(\rho=12)\cdot C(\boldsymbol{\pi_{1}})$ (see [36]), we have directly from induction that $C(\boldsymbol{\pi})=r(\mathcal{T}(\boldsymbol{\pi}))+1$. We summarize this as:
Lemma 4.
Let $\boldsymbol{\pi}\sim\operatorname{Unif}(\operatorname{B}_{n})$. Then $C(\boldsymbol{\pi})=r(\mathcal{T}(\boldsymbol{\pi}))+1\sim\ell(\mathcal{T}(\boldsymbol{\pi}))+1$.
Remark 5.
In both the simple and nonsimple cases, different structural features of the permutation correspond directly to the same geometric quantity in the associated butterfly tree. For simple butterfly permutations—constructed via Kronecker products—the right edge length satisfies $r(\mathcal{T}(\pi))+1=\operatorname{LDS}(\pi)$, linking it to the length of the longest decreasing subsequence (cf. Lemma 3). In contrast, for nonsimple butterfly permutations—formed using wreath products—we find that now $r(\mathcal{T}(\pi))+1=C(\pi)$, aligning the right edge instead with the number of cycles.
Hence, as seen also in [36], then the moments of $L_{n}$ and $R_{n}$ can be computed exactly (cf. [36, Proposition 9] as well as a full asymptotic distributional limit (cf. [36, Theorem 10]). In particular, a direct computation yields
$\mathbb{E}L_{n}$ $=\mathbb{E}R_{n}=\left(\frac{3}{2}\right)^{n}-1$ (17)
$\mathbb{E}L_{n}^{2}$ $=\mathbb{E}R_{n}^{2}=\frac{4}{3}\left(\frac{3}{2}\right)^{2n}-\frac{7}{3}\left(\frac{3}{2}\right)^{n}+1$ (18)
We now have sufficient tools to commence proving Theorem 4.
Since $L_{n}\sim R_{n}$, then from (16) and using $\max(x,y)=(x+y+|x-y|)/2$, we have
$\mathbb{E}H_{n+1}=\frac{1}{2}\mathbb{E}(L_{n}^{\prime}+1+H_{n}+H_{n}^{\prime}+|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|)=\frac{1}{2}\left(\frac{3}{2}\right)^{n}+\mathbb{E}H_{n}+\frac{1}{2}\mathbb{E}|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|.$
Using $\mathbb{E}|Y|\geq|\mathbb{E}Y|$ and (17), we have
$\mathbb{E}|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|\geq|\mathbb{E}L_{n}^{\prime}+1+\mathbb{E}H_{n}-\mathbb{E}H_{n}^{\prime}|=|\mathbb{E}L_{n}^{\prime}+1|=\left(\frac{3}{2}\right)^{n}.$
This yields further then the lower bound
$\mathbb{E}H_{n+1}\geq\left(\frac{3}{2}\right)^{n}+\mathbb{E}H_{n},$
which simplifies to
$\mathbb{E}H_{n}\geq\sum_{k=0}^{n-1}\left(\frac{3}{2}\right)^{k}=2\left(\frac{3}{2}\right)^{n}-2.$ (19)
Remark 6.
This lower bound on the mean height of uniform nonsimple butterfly trees matches exactly the mean height for uniform simple butterfly trees (cf. Theorem 3). A similar phenomenon appears for the LIS, where the average LIS under uniform simple butterfly permutations serves as a lower bound for the nonsimple case (cf. Theorems 7 and 8).
For an upper bound, we have
$\mathbb{E}|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|\leq\mathbb{E}|L_{n}^{\prime}+1|+\mathbb{E}|H_{n}-H_{n}^{\prime}|\leq\left(\frac{3}{2}\right)^{n}+\sqrt{2\operatorname{Var}(H_{n})}.$
via the triangle inequality, Cauchy-Schwarz, and (17). Hence,
$\mathbb{E}H_{n+1}\leq\left(\frac{3}{2}\right)^{n}+\mathbb{E}H_{n}+\frac{1}{2}\sqrt{2\operatorname{Var}(H_{n})}$
using (16). To work with $\operatorname{Var}(H_{n})$, we can first tackle $\operatorname{Var}(L_{n})$. Using (18), then
$\mathbb{E}H_{n+1}^{2}$ $=\frac{1}{2}\mathbb{E}\max((L_{n}^{\prime}+1+H_{n})^{2},H_{n}^{\prime 2})+\frac{1}{2}\mathbb{E}\max((R_{n}+1+H_{n}^{\prime})^{2},H_{n}^{2})$
$=\mathbb{E}\max((L_{n}^{\prime}+1+H_{n})^{2},H_{n}^{\prime 2})$
$=\frac{1}{2}\mathbb{E}[(L_{n}^{\prime}+1+H_{n})^{2}+H_{n}^{\prime 2}+(L_{n}^{\prime}+1+H_{n}+H_{n}^{\prime})|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|]$
$=\frac{1}{2}\mathbb{E}[(L_{n}^{\prime}+1)^{2}+H_{n}^{2}+2H_{n}+2L_{n}^{\prime}H_{n}+H_{n}^{\prime 2}+(L_{n}^{\prime}+1+H_{n}+H_{n}^{\prime})|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|]$
$=\frac{1}{2}\mathbb{E}[(L_{n}+1)^{2}]+\mathbb{E}H_{n}^{2}+(\mathbb{E}L_{n}+1)(\mathbb{E}H_{n})+\frac{1}{2}\mathbb{E}[(L_{n}^{\prime}+1+H_{n}+H_{n}^{\prime})|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|].$
Now note
$(\mathbb{E}H_{n+1})^{2}$ $=\left(\frac{1}{2}(\mathbb{E}L_{n}+1)+\mathbb{E}H_{n}+\frac{1}{2}\mathbb{E}|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|\right)^{2}$
$=\frac{1}{4}(\mathbb{E}L_{n}+1)^{2}+(\mathbb{E}H_{n})^{2}+\frac{1}{4}(\mathbb{E}|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|)^{2}+(\mathbb{E}L_{n}+1)(\mathbb{E}H_{n})$
$\hskip 24.0pt+(\mathbb{E}H_{n})(\mathbb{E}|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|)+\frac{1}{4}(\mathbb{E}L_{n}+1)(\mathbb{E}|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|).$
It follows
$\operatorname{Var}(H_{n+1})$ $=\mathbb{E}H_{n+1}^{2}-(\mathbb{E}H_{n+1})^{2}$
$=\frac{1}{4}\mathbb{E}[(L_{n}+1)^{2}]+\frac{1}{4}\operatorname{Var}(L_{n})+\operatorname{Var}(H_{n})+\frac{1}{2}\mathbb{E}[(L_{n}^{\prime}+1+H_{n}+H_{n}^{\prime})|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|]$
$\hskip 24.0pt-\frac{1}{4}(\mathbb{E}L_{n}+2+4\mathbb{E}H_{n})(\mathbb{E}|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|).$
Now note
$\mathbb{E}[(L_{n}^{\prime}+1+H_{n}+H_{n}^{\prime})|L_{n}^{\prime}+1+H_{n}-H_{n}^{\prime}|]$
$\hskip 24.0pt\leq\mathbb{E}[(L_{n}^{\prime}+1+H_{n}+H_{n}^{\prime})(L_{n}^{\prime}+1+|H_{n}-H_{n}^{\prime}|)]$
$\hskip 24.0pt=\mathbb{E}[(L_{n}^{\prime}+1)^{2}+H_{n}L_{n}^{\prime}+H_{n}+H_{n}^{\prime}L_{n}^{\prime}+H_{n}^{\prime}+H_{n}|H_{n}-H_{n}^{\prime}|+H_{n}^{\prime}|H_{n}-H_{n}^{\prime}|]$
$\hskip 60.0pt+\mathbb{E}[L_{n}^{\prime}|H_{n}-H_{n}^{\prime}|]+\mathbb{E}[|H_{n}-H_{n}^{\prime}|]$
$\hskip 24.0pt=\mathbb{E}[(L_{n}+1)^{2}]+(\mathbb{E}H_{n})(\mathbb{E}L_{n})+2\mathbb{E}H_{n}+\mathbb{E}[H_{n}L_{n}]+2\mathbb{E}[H_{n}|H_{n}-H_{n}^{\prime}|]$
$\hskip 60.0pt+\mathbb{E}[L_{n}|H_{n}-H_{n}^{\prime}|]+\mathbb{E}[|H_{n}-H_{n}^{\prime}|]$
$\hskip 24.0pt\leq\mathbb{E}[(L_{n}+1)^{2}]+(\mathbb{E}H_{n})(\mathbb{E}L_{n}+2)+\sqrt{(\operatorname{Var}(H_{n})+(\mathbb{E}H_{n})^{2})(\mathbb{E}L_{n}^{2})}$
$\hskip 60.0pt+2\sqrt{2(\operatorname{Var}(H_{n})+(\mathbb{E}H_{n})^{2})\operatorname{Var}(H_{n})}+\sqrt{2\mathbb{E}L_{n}^{2}\operatorname{Var}(H_{n})}+\sqrt{2\operatorname{Var}(H_{n})}$
using the triangle and Cauchy-Schwarz inequalities. Hence, we have
$\operatorname{Var}(H_{n+1})$ $\leq\operatorname{Var}(H_{n})+\frac{1}{4}(4\mathbb{E}L_{n}^{2}+4(\mathbb{E}L_{n}+1)-(\mathbb{E}L_{n}+1)^{2})+\frac{1}{2}(\mathbb{E}H_{n})(\mathbb{E}L_{n}+2)$
$\hskip 24.0pt+\frac{1}{2}\sqrt{(\operatorname{Var}(H_{n})+(\mathbb{E}H_{n})^{2})(\mathbb{E}L_{n}^{2})}+\sqrt{2(\operatorname{Var}(H_{n})+(\mathbb{E}H_{n})^{2})\operatorname{Var}(H_{n})}$
$\hskip 24.0pt+\frac{1}{2}\sqrt{2\mathbb{E}L_{n}^{2}\operatorname{Var}(H_{n})}+\frac{1}{2}\sqrt{2\operatorname{Var}(H_{n})}.$
Let $\lambda=3/2$. Note then obviously $\lambda^{n}-1=\mathbb{E}L_{n}\leq\lambda^{n}$ while $\mathbb{E}L_{n}^{2}=\frac{4}{3}\lambda^{2n}-(\frac{7}{3}\lambda^{n}-1)\leq\frac{4}{3}\lambda^{2n}$. It follows
$\operatorname{Var}(H_{n+1})$ $\leq\operatorname{Var}(H_{n})+\frac{1}{4}\left(4\cdot\frac{4}{3}\lambda^{2n}+4\lambda^{n}-\lambda^{2n}\right)+\frac{1}{2}(\mathbb{E}H_{n})(\lambda^{n}+1)$
$\hskip 24.0pt+\frac{1}{2}\sqrt{\frac{4}{3}\lambda^{2n}}\sqrt{\operatorname{Var}(H_{n})+(\mathbb{E}H_{n})^{2}}+\sqrt{2(\operatorname{Var}(H_{n})+(\mathbb{E}H_{n})^{2})\operatorname{Var}(H_{n})}$
$\hskip 24.0pt+\frac{1}{2}\sqrt{\frac{8}{3}\lambda^{2n}}\sqrt{\operatorname{Var}(H_{n})}+\frac{1}{2}\sqrt{2\operatorname{Var}(H_{n})}$
$\leq\operatorname{Var}(H_{n})+\frac{13}{12}\lambda^{2n}+\lambda^{n}+\frac{1}{2}(\mathbb{E}H_{n})(\lambda^{n}+1)$
$\hskip 24.0pt+\sqrt{\frac{1}{3}}\lambda^{n}\sqrt{\operatorname{Var}(H_{n})+(\mathbb{E}H_{n})^{2}}+\sqrt{2}\cdot\sqrt{(\operatorname{Var}(H_{n})+(\mathbb{E}H_{n})^{2})\operatorname{Var}(H_{n})}$
$\hskip 24.0pt+\sqrt{\frac{2}{3}}\lambda^{n}\sqrt{\operatorname{Var}(H_{n})}+\frac{1}{2}\sqrt{2\operatorname{Var}(H_{n})}.$
Let $a_{n},b_{n}\geq 0$ with $a_{0}=b_{0}=0$ such that
$a_{n+1}$ $=\lambda^{n}+a_{n}+\frac{1}{2}\sqrt{2b_{n}}$
$b_{n+1}$ $=b_{n}+\frac{13}{12}\lambda^{2n}+\lambda^{n}+\frac{1}{2}(\lambda^{n}+1)a_{n}+\sqrt{\frac{1}{3}}\lambda^{n}\sqrt{b_{n}+a_{n}^{2}}$
$\hskip 24.0pt+\sqrt{(b_{n}+a_{n}^{2})2b_{n}}+\left(\sqrt{\frac{1}{3}}\lambda^{n}+{\frac{1}{2}}\right)\sqrt{2b_{n}}.$
Our goal is to bound both $\mathbb{E}H_{n}$ and $\operatorname{Var}(H_{n})$ in terms of $a_{n},b_{n}$. By construction and induction, we have
$\mathbb{E}H_{n}\leq a_{n}\ \mbox{and}\ \operatorname{Var}(H_{n})\leq b_{n}\qquad\mbox{for all }n.$
Next, observe that
$a_{n+1}^{2}=\lambda^{2n}+a_{n}^{2}+\frac{1}{2}b_{n}+2\lambda^{n}a_{n}+(\lambda^{n}+a_{n})\sqrt{2b_{n}}.$
We aim to find a constant $C\geq 0$ such that $b_{n}\leq Ca_{n}^{2}$. This allows us to bound
$a_{n+1}\leq\lambda^{n}+\xi a_{n},$
which has general upper bound recursion solution
$a_{n}\leq A\lambda^{n}+B\xi^{n},$
for $\xi=1+\sqrt{2C}/2$. Moreover, if $\frac{1}{2}<C<2$ then $\lambda<\xi<2$, which would thus yield a sub-linear upper bound on $a_{n}$ and hence $\mathbb{E}H_{n}$, with exponent $\log_{2}\xi<1$.
The remainder of this section will verify that this bound holds with the choice
$C^{*}:=1+\sqrt{8\sqrt{2}-11}\approx 1.5601.$ (20)
This in turn yields the desired sublinear upper bound on $\mathbb{E}h(\mathcal{T}(\pi_{n}))$ for $\pi_{n}\sim\operatorname{Unif}(\operatorname{B}_{n})$, as stated in Theorem 4, with exponent $\beta=\log_{2}\xi\approx 0.913189$.
Lemma 5.
Let $C\geq C^{*}$. Then $b_{n}\leq Ca_{n}^{2}$ for all $n\geq 0$.
Proof.
Now this trivially holds for any $C$ when $n=0$, while by above we have $a_{1}=\frac{3}{2}$ and $b_{1}=\frac{13}{12}+1=\frac{25}{12}$, so this holds as long as $C\geq\frac{25}{27}$. Now assume the result holds for $n\geq 1$. Recall by above $a_{n}\geq\mathbb{E}H_{n}\geq 2\lambda^{n}-2$, so $\lambda^{n}\leq\frac{1}{2}a_{n}+1\leq\frac{3}{2}a_{n}$ for all $n\geq 1$. Using the triangle inequality, we have $\sqrt{b_{n}+a_{n}^{2}}\leq\sqrt{b_{n}}+a_{n}$, and so
$b_{n+1}$ $\leq(1+\sqrt{2})b_{n}+\frac{13}{12}\lambda^{2n}+\left(\left(\frac{1}{2}+\sqrt{\frac{1}{3}}\right)\lambda^{n}+2\right)a_{n}$
$\hskip 24.0pt+\left(\left(\sqrt{\frac{1}{6}}+\sqrt{\frac{1}{3}}\right)\lambda^{n}+a_{n}+{\frac{1}{2}}\right)\sqrt{2b_{n}}.$
To get a potential value of $C$, we next compute
$b_{n+1}-Ca_{n+1}^{2}$ $\leq-2\left(\left(C-\frac{1}{2}\left(\frac{1}{2}+\sqrt{\frac{1}{3}}\right)\right)\lambda^{n}-1\right)a_{n}-\left(\left(C-\left(\sqrt{\frac{1}{6}}+\sqrt{\frac{1}{3}}\right)\right)\lambda^{n}-\frac{1}{2}\right)\sqrt{2b_{n}}$
$\hskip 24.0pt-\left(C-\frac{13}{12}\right)\lambda^{2n}+\left[\left(1+\sqrt{2}-\frac{C}{2}\right)b_{n}-Ca_{n}^{2}-(C-1)a_{n}\sqrt{2b_{n}}\right]$
[…]
One way to ensure $b_{n+1}\leq Ca_{n+1}^{2}$ is if $C$ is chosen such that the upper bound has non-positive coefficients for each of the three terms using $a_{n}$, $\sqrt{2b_{n}}$ and $\lambda^{2n}$, i.e., I through III, along with ensuring the term inside the square brackets IV is also non-positive. This will be our approach here. Future refinements can allow some positive values across the four terms as long as the overall sum remains non-positive.
Note $\frac{1}{2}(\frac{1}{2}+\sqrt{\frac{1}{3}})+1\approx 1.5387$, $\sqrt{\frac{1}{6}}+\sqrt{\frac{1}{3}}+\frac{1}{2}\approx 1.48560$, and $\frac{13}{12}\approx 1.08333$ yield lower bounds on potential values of $C$ so that each respective term I through III is non-positive for all $n$. Hence, ensuring $C$ exceeds the maximum $\frac{1}{2}(\frac{1}{2}+\sqrt{\frac{1}{3}})+1$ then serves as a sufficient value to make sure all of these three terms are non-positive.
To further refine our choice of $C$, we can find the best bound to ensure the remaining term IV is non-positive. Using the inductive hypothesis for $b_{n}\leq Ca_{n}^{2}$ for all $n$, then
$b_{n}\leq\sqrt{C}a_{n}\sqrt{b_{n}}\leq Ca_{n}^{2}.$
It follows then the remaining term satisfies the bound
$\framebox{IV}=\left(1+\sqrt{2}-\frac{C}{2}\right)b_{n}-Ca_{n}^{2}-(C-1)a_{n}\sqrt{2b_{n}}\leq f(C)\cdot b_{n}$
for
$f(x)=\sqrt{2}-\frac{x}{2}-\sqrt{\frac{2}{x}}\cdot(x-1).$
We now show $f(x)\leq 0$ iff $x\geq C^{*}$ for $C^{*}\approx 1.5601$ from (20). This follows since $f^{\prime}(x)<0$ for $x>0$ with $f(x)\to+\infty$ as $x\to 0+$ and $f(x)\to-\infty$ as $x\to\infty$ yields $f$ has a unique 0, which can be directly verified to be $C^{*}$ using Cardano’s formula to keep the (unique) positive root of the associated cubic polynomial $yf(y^{2})$ in $y=\sqrt{x}$ (and uniqueness for such a positive root follows from Descartes rule of signs). Hence, IV is non-positive if $C\geq C^{*}$. Since $1.5601\approx C^{*}>\frac{1}{2}(\frac{1}{2}+\sqrt{\frac{1}{3}})\approx 1.5387$, then this choice also ensures I through III are each non-positive. Applying this to (21) then yields $b_{n+1}\leq Ca_{n+1}^{2}$. This completes the inductive step.
Lemma 5 now establishes $b_{n}\leq Ca_{n}^{2}$ for any $C\geq C^{*}$. In particular, this holds for $C=C^{*}$, which we will now assume to be the case going forward. Now we have
$a_{n+1}=\lambda^{n}+a_{n}+\frac{1}{2}\sqrt{2b_{n}}\leq\lambda^{n}+\xi a_{n},\qquad\xi=1+\frac{\sqrt{2C^{*}}}{2}=\frac{1}{2}\left(1+\sqrt{2}+\sqrt{2\sqrt{2}-1}\right)\approx 1.88320.$
Solving the upper bound recursion $a_{n+1}\leq\lambda^{n}+\xi a_{n}$ with $a_{0}=0$ yields the bound
$\mathbb{E}H_{n}\leq a_{n}\leq\frac{\xi^{n}-\lambda^{n}}{\xi-\lambda}.$ (22)
Hence, noting $1.51.88320\approx\xi>\lambda=$ 1.5, we have thus the desired sub-linear upper bound
$\mathbb{E}H_{n}\leq dN^{\beta}\cdot(1+o(1)),\qquad\beta=\log_{2}\xi\approx 0.913189,\qquad\quad d=\frac{1}{\xi-\lambda}=\frac{2}{\sqrt{2C^{*}}-1}\approx 2.60958.$
This completes the proof of Theorem 4.
Figure 7: Comparison of $N-1$ to the lower and upper bounds from Theorem 4
Remark 7.
[…]
Remark 8.
[…]